### v1

# Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases

Abstract.
In recent years, there has been a growing recognition of the need to incorporate lay-people’s input into the governance and acceptability assessment of AI usage. However, how and why people judge different AI use cases to be acceptable or unacceptable remains under-explored. In this work, we investigate the attitudes and reasons that influence people’s judgments about AI’s development via a survey administered to demographically diverse participants (N=197). We focus on ten distinct professional (e.g., Lawyer AI) and personal (e.g., Digital Medical Advice AI) AI use cases to understand how characteristics of the use cases and the participants’ demographics affect acceptability. We explore the relationships between participants’ judgments and their rationales such as reasoning approaches (cost-benefit reasoning vs. rule-based). Our empirical findings reveal number of factors that influence acceptance such as general negative acceptance and higher disagreement of professional usage over personal, significant influence of demographics factors such as gender, employment, and education as well as AI literacy level, and reasoning patterns such as rule-based reasoning being used more when use case is unacceptable. Based on these findings, we discuss the key implications for soliciting acceptability and reasoning of AI use cases to collaboratively build consensus. Finally, we shed light on how future FAccT researchers and practitioners can better incorporate diverse perspectives from lay people to better develop AI that aligns with public expectations and needs.
1. Introduction
There are growing calls to regulate AI’s development and integration into society (Pistilli et al., 2023). These efforts, as reflected in the EU AI Act (Parliament, 2023), NIST AI Risk Management framework (of Standards and Technology, 2023), and recent U.S. Executive Order (exe, 2023), have resulted in discussions about whether certain AI use cases should be pursued at all. Despite much progress in this area, it is still not clear how to determine which use cases should be pursued or more heavily regulated. Further, little is known about how lay community members, especially those from marginalized groups, feel about the development of specific AI use cases (ada, 2023; Suresh et al., 2024).
One significant challenge when evaluating the acceptability and impact of specific AI use cases111By use cases, we mean specific scenarios, applications, or problems that an AI system is designed to solve or assist within real-world contexts. is that there can be both positive and negative effects, depending on the context(Mun et al., 2024). For instance, while educational AI can provide affordable and accessible personal tutor, it can also lead to over-reliance of students and diminish the goal of education (Times, 2024; Zhai et al., 2024). To develop a generalizable approach to making decisions about AI use cases, ideally we would understand how people resolve these conflicts. More specifically, first, it is essential to understand differences in judgments about acceptability and likely usage across use cases, and how such judgments relate to scenario characteristics. Second, we need better understanding of the personal factors influencing these judgments, especially as they relate to demographic differences (Kingsley et al., 2024). Third, we need to better understand the reasoning strategies participants use when making judgments about AI use cases, and how those strategies do or do not relate to the judgments that are ultimately made. To form governance and policy decisions that anticipate and address disagreements about the development or regulation about specific AI use cases, these understandings are crucial, especially across groups of people with diverse backgrounds, experiences, and familiarity with AI.
To address these needs, in this work we examine how and why lay people judge various AI use cases as acceptable or unacceptable, asking the following research questions:
RQ1
How do judgments of acceptability vary across a set of distinct AI use cases and their characteristics?
RQ2
What attributes or characteristics of people explain the variation in acceptability judgments?
RQ3
How do people reason through acceptability judgments of AI use cases?
To answer these questions, we develop a survey to collect judgments and reasoning processes of 197 demographically diverse participants with varying levels of experience with AI. We ask participants to report whether a certain AI use case should be developed or not, whether they would use such a system, and ask them to provide rationales for their judgment and conditions that would cause them to change their judgments (Figure 1). We examine ten different AI use cases222We focus on text-based, non-embodied, digital systems, and while we do not specifically discuss the AI user and subject, in our use case description, we follow three of the five concepts used in EU AI Act to describe high risk use cases (Golpayegani et al., 2023): the domain, purpose, and capabilities.. To account for differences between sectors or domains, we select use cases in two categories, professional and personal use, and vary them by required entry-level education and EU AI risk level (Table 1).
To meaningfully differentiate and analyze participants’ reasons and reasoning strategies, we borrow concepts from moral psychology and philosophy. We investigate participants’ rationales through two distinct but sometimes overlapping reasoning patterns: cost-benefit reasoning, which assesses expected outcomes (e.g., ”using AI for this task would save time”), and rule-based reasoning, which evaluates the intrinsic values of the action itself (e.g., ”having humans perform this task would be inherently wrong”) (Cushman, 2013; Cheung et al., 2024). We further explore the moral foundations reflected in participants’ reasoning, with moral foundations theory333We used the five foundational dimensions: Care, Fairness, Loyalty, Authority, and Purity. Although these dimensions have been updated to encompass a broader range of values beyond WEIRD (White, Educated, Industrialized, Rich, and Democratic) populations (Atari et al., 2023), we selected this version for survey brevity. (Graham et al., 2011, 2008). Additionally, to understand aspects of AI that raise concerns, we employ three dimensions based on prior studies (Solaiman et al., 2023; Mun et al., 2024): functionality (system capabilities like performance, bias, and privacy), usage (context of system integration, such as supervision, misuse, or unintended use), and societal impact (effects on individuals, communities, and society, such as job loss and over-reliance).
Our empirical results show general higher acceptance of personal use cases over professional. While both categories of use cases show decreased acceptance with increased entry level education and risk, professional use cases display more variability and disagreements across judgments (RQ1). Acceptability significantly varied among demographic groups and levels of AI literacy, with lower acceptability observed particularly among non-male participants and those familiar with AI ethics (RQ2). Finally, our results show varying distribution of reasoning types across acceptability decisions with rule-based reasoning being associated with negative acceptance as well as concern for societal impact. Further qualitative analysis reveal rules such as the need for humanness in certain use cases whether it be for empathy or interaction (RQ3).
Our findings shed novel light onto the diversity of people’s acceptability and reasoning of AI uses in distinct domains and risk levels. We conclude with a discussion highlighting three key implications: first, diverse methodologies are needed to effectively analyze use cases and their characteristics; second, involving diverse stakeholders is crucial for assessing the acceptability of AI applications, particularly in workplaces; and third, further investigation into human reasoning processes about AI, notably rule-based reasoning, is needed to inform consensus-building in policy making.
2. Related Works
In this section, we briefly summarize the background and related work towards assessing acceptability and impact of AI use cases. In each subsection, we highlight how our work extends prior work.
2.1. Assessing Impact of AI
Recent years have witnessed increasing calls from academics (Kieslich et al., 2023; Hecht et al., 2021; Bernstein et al., 2021; Ashurst et al., 2020; Lin et al., 2020; CVPR, Ethics Guidelines, 2023; ACL, Ethic Policy, 2023; ICML, Publication Ethics, 2023; Olteanu et al., 2023; Center for Advanced Study in the Behavioral Sciences, 2020), government (The Ada Lovelace Insitute, 2022; The White House, 2023; National Artificial Intelligence Research Resource Task Force, 2023), civil society (on AI, 2021; Ada Lovelace Institute, 2022; The Ada Lovelace Insitute, 2022; Society, 2023; Metcalf et al., 2021; Reisman et al., 2018), and industry (Microsoft, 2022a, b; Walker and Croak, 2021; OpenAI, 2022; Hecht et al., 2021; Deng et al., 2024) to assess the impact of AI systems designed and developed by AI researchers and practitioners. This effort has particularly highlighted the need to understand the positive impact while grappling with the potential negative impact of integrating AI into certain products and services that affect people’s daily lives (Ashurst et al., 2020; Lin et al., 2020; Hecht et al., 2021). In response, researchers in FAccT, HCI, and AI have developed tools and processes to support AI researchers and practitioners in anticipating the impact of AI systems they developed(Wang et al., 2024; Kieslich et al., 2023; Buçinca et al., 2023; Deng et al., 2024; Weidinger et al., 2022). For example, many have developed AI impact taxonomies or checklists to help developers categorize AI impact(Weidinger et al., 2022; Microsoft, 2022b). Wang et al. and Deng et al. developed tools and templates to support industry AI developers and researchers in assessing the potential negative societal impact of their work, such as job displacement or stereotyping social groups.
However, this prior work primarily focuses on supporting AI experts rather than diverse lay people’s impact assessments of potential AI use cases. Our work extends these prior efforts by understanding diverse (and sometimes conflicting) perspectives on both positive and negative impact of AI use cases from lay people, as a crucial step to complement the AI impact assessments conducted by AI researchers and practitioners.
2.2. Understanding People’s Perceptions of AI Use Cases
Responding to the calls on meaningfully engaging lay people in assessing the impact of specific AI use cases, prior work have started to understand lay people’s perceptions of AI use case (Buçinca et al., 2023; Kieslich et al., 2023; Mun et al., 2024; Kingsley et al., 2024). Among other findings, this prior work revealed a substantial amount of disagreement regarding the desired behavior of AI, primarily due to the subjectivity inherent in certain tasks (e.g., toxicity detection (Sap et al., 2019; Blodgett et al., 2020), image captioning (Zhao et al., 2021)) and ambiguous ethical implication of decisions made by AI for certain tasks (e.g., self driving cars (Awad et al., 2018), medical AI (Chen et al., 2023), predictive analysis (Barocas and Selbst, 2016)). Work done by Mun et al. highlighted that lay people can envision diverse set of harms specific to different AI use cases, complementary to those defined by experts. Another line of work also begins to examine how factors such as demographic backgrounds and previous exposure to discrimination can affect people’s sensitivity towards potential AI harms (Kingsley et al., 2024).
Our work extends this prior work by examined the detailed reasoning processes of lay people regarding the acceptability and the trade-offs between positive and negative impacts of AI use cases. In particular, we draw on model decision theory framework, such as the moral foundations developed by Graham et al., to design a survey flow (See Figure 1 to solicit lay people’s decision-making processes (and potential moral conflicts) when assessing both the benefits and harms of concrete AI use cases.
2.3. Background: Moral Decision Making
Morality, characterized by diverse values across cultures and social groups, aims to suppress selfishness to facilitate social life (Kesebir and Haidt, 2010). To understand decision-making in AI use cases, we draw on moral psychology and dual system theory. We examine two decision-making systems: cost-benefit reasoning, which assesses outcomes and consequences, and rule-based reasoning, focusing on norms, rules, and virtues (Cushman, 2013; Cheung et al., 2024). These correspond to utilitarian reasoning (maximizing good) and deontological reasoning (duties and rights), respectively. Additionally, we apply moral foundations theory (Graham et al., 2008) to identify values and potential moral conflicts in AI development.
3. Study Design and Data Collection
To understand the decision making factors and reasoning patterns of a diverse population regarding AI use cases, we conducted a survey-based study with demographically diverse participants. In this section, we discuss the selection of use cases (§ 3.1), survey design (§ 3.2), and data collection details and participant demographics (§ 3.3).
3.1. Use Cases
To understand how different characteristics of AI use cases can impact judgments and decision making processes, we crafted ten different AI use cases. We first chose two broad application categories frequently mentioned by the public in previous works (Kieslich et al., 2024; Mun et al., 2024): AI in personal, everyday usage where participants could uniformly consider themselves as AI users and AI in professional usage where AI takes on a role thus far done by a human as a profession, e.g., Lawyer AI. We then developed five use cases in the AI in professional usage category that varied according to the level of education required for entry and five use cases in AI in personal category that varied according to the risk level assigned by the EU AI act. All the personal use cases were in the health domain, reflecting a key domain of interest conveyed by lay users in previous works (Mun et al., 2024; Kieslich et al., 2024).
| Use Case | Factor | Description |
| Professional Use Cases | ||
| Lawyer | Doctoral/Professional Degree | Advises clients on digital legal proceedings/transactions. |
| Elementary School Teacher | Bachelor’s degree | Teaches academic skills at the elementary school level. |
| IT Support Specialist | Some college, no degree | Maintains computer networks and provides technical help. |
| Government Eligibility Interviewer | High school diploma | Determine eligibility for government programs/resources. |
| Telemarketer | No formal education | Solicits donations or orders over the telephone. |
| Personal Use Cases | ||
| Digital Medical Advice | High Risk | Provide medical assessments prior to medical consultations. |
| Customized Lifestyle Coach | High / Limited Risk | Personalized advice for healthy living and wellness. |
| Personal Health Research | Limited Risk | Summarizes research related to personal health issues. |
| Nutrition Optimizer | Limited / Low Risk | Personalize meals and optimize nutritional intake. |
| Flavorful Swaps | Low Risk | Suggest delicious and healthy alternatives food options. |
Professional Use Case Scenarios
For the first area of focus, AI in labor replacement, we collected jobs listed in the U.S. census bureau444https://www.bls.gov/ooh/occupation-finder.htm and sorted them according to entry level education required as stated in the census. We chose education level as it has been tightly linked to socioeconomic and occupational status (Svensson, 2006; Evetts, 2006). We selected jobs that have a large portion of digital or intellectual components with minimal requirement for embodiment resulting in following five professional roles: Lawyer, Elementary school teacher, IT support specialist, Government support eligibility interviewer, and Telemarketer. See Table 1 for further details.
Personal Use Case Scenarios
To understand the acceptability of different health applications in personal and private life, we adapted the descriptions of personal use cases written by participants from prior works (Mun et al., 2024; Kieslich et al., 2024). The research team iteratively refined the descriptions to reflect risk levels according to EU AI Act, and confirmed agreement with categories assigned by GPT-4, following Herdel et al.. See Table 1 for further details.
3.2. Survey Design
Our survey presents participants with five use case descriptions, all from either the professional or personal category (randomly assigned and presented in random order) (see § 3.1 for details). After each description, participants answer: ”Do you think a technology like this should be developed?” (Q1) and then, ”How confident are you in your above answer?” (Q2). They provide open-text rationales by completing the prompt, ”[Use Case] should be developed because…” (Q3), tailored to their Q1 response (”should” for ”Yes” and ”should not” for ”No”). Participants also specify a condition for switching their decision with, ”[Use Case] should not be developed if…” (Q4), adjusted similarly to Q1. Subsequently, they answer, ”If [Use Case] existed, would you use its service?” (Q5) and express confidence with, ”How confident are you in your above answer?” (Q6). Refer to Table 6 in the Appendix for the exact wording of the questions.
Collecting Participant Characteristics
Following the main survey, we asked participants questions about their AI literacy and demographics to explore various factors affecting perception of AI acceptance. We adopted a shortened version of AI literacy questionnaires from previous works (Wang et al., 2023; Mun et al., 2024) with four AI literacy aspects, AI awareness, usage, evaluation, and ethics, and two additional questions for generative AI, usage frequency and familiarity with limitations. We collected demographic information of the participants such as race, gender, age, sexual orientation, religion, employment status, income, and level of education. Additionally, we collected information about chronicity, i.e., prolonged experiences of everyday discrimination, of their discrimination experiences (if any) following Kingsley et al..
3.3. Data Collection and Participant Demographics
We used Prolific555https://www.prolific.com to recruit participants. To represent diverse sample, we stratified our recruitment by the ethnicity categories (White, Mixed, Asian, Black and Other) and age (18-48, 49-100) provided by Prolific. We also added criteria for quality such as survey approval rating and number of previous surveys completed. Our study was approved by IRB at our institutions, and we paid 12 USD/hour. Our final sample consisted of 197 participants across two categories, with professional usage assigned to 100 participants and personal to 97. See Appendix A.2 for further details on participants.
4. Analysis Methods
Our surveys consisted of both multiple choice (numerical) and open-text questions designed to answer our research questions. In this section, we detail our process for numerical (§ 4.1) and open-text (§ 4.2) analysis.
4.1. Multiple Choice Analysis
We analyzed the judgment and confidence ratings by mapping judgment (Q1, Q5) to 1 (“Should be developed”, “Would use”) or -1 (“Should not be developed”, “Would not use”) and confidence (Q2, Q6) to a scale from 1 to 5. We used numerically converted judgment, confidence, and combined (judgmentconfidence; -5 to 5) values as dependent variables in our analysis. We used repeated-measures ANOVAs to understand the differences in mean responses between conditions/groups and linear mixed effects regression models (lmer) to better understand the effects of specific factors. We included a subject-specific random effect when using ANOVA and regression models and added a use-case-specific random effect when applicable. We factorized demographic responses for analysis with the exception of discrimination chronicity, which we aggregated to a numerical value (Kingsley et al., 2024; Michaels et al., 2019). We also converted responses to AI literacy questions to numerical values for analysis.
4.2. Open-response Analysis
To assess the reasoning methods used by the participants, we analyzed the open-text responses on elaborations to their decisions (Q3) and circumstances in which their decisions would switch (Q4) along the following three dimensions: reasoning types (cost-benefit, rule-based, both, unclear), reference to moral foundations (Care, Fairness, Purity, Authority, Loyalty), and switching conditions (Functionality, Usage, Societal Impact). By analyzing reasoning types and moral values reflected in the participants’ justifications, we aim to characterize how participants made their decisions, and by analyzing various factors such as primary concerns and stakeholders, we aim to discover what aspects were salient for the participants in their decisions.
Classification and Aggregation
We classified participants’ responses to Q3 (elaboration of judgment) and Q4 (conditions for switching decisions) using OpenAI’s o1-mini666o1-mini-2024-09-12. To validate the model’s classification performance, results were compared with a reference set of 100 samples annotated by three independent annotators, comprised of team members and a professional annotator. Initially, each annotator independently assessed the data, and then consensus was reached through discussion to establish a gold standard set. The inter-rater agreement between the gold standard and o1-mini’s annotations was evaluated using Gwet’s AC1 metric, chosen for its robustness with infrequent labels (Wongpakaran et al., 2013). The agreement levels varied, ranging from moderate to almost perfect with no dimension below moderate. Annotations for cost-benefit reasoning, rule-based reasoning, and authority reached near-perfect agreement; purity and usage achieved substantial agreement; the rest showed moderate agreement. Due to a lack of sufficient test samples and minimal occurrences in the annotated data, the moral value dimension Loyalty was excluded from further analysis. The annotations were converted into a one-hot encoding format for statistical analysis. See Appendix B for further details.
5. Findings
Our work aims to uncover variations in acceptability of AI use cases and factors and reasoning processes that underlie these judgments. In this section, we discuss our findings about the judgments of the AI use cases (§5.1), personal factors that may influence the decision such as demographics and AI literacy (§5.2), and factors in rationales that could uncover reasoning processes that lead to judgments (§5.3).
5.1. RQ1. Perceptions and Disagreements of Use Cases
5.1.1. Use Case Factors
In our analysis, we investigated the effects of use cases on participants’ judgments of two categories—professional and personal—along with ten specific use cases on judgments and confidence about their development and usage. Figure 2 illustrates significant differences in acceptability depending on the categories. Notably, personal use cases had higher acceptability (; ) than professional use cases (; ). These differences are more pronounced when judgments are weighted by confidence. Although levels of confidence did not differ significantly between categories, they were fairly high for both categories, and personal category exhibited slightly lower confidence for usage.
Professional Use Cases
Exploring specific use cases within the professional category, we observed that Elementary School Teacher AI (; ) had the lowest acceptability for both types of judgments followed by Lawyer AI () and Telemarketer AI (). Interestingly, IT Support Specialist AI had the highest acceptability (; ). While all other use cases showed higher acceptability for usage over development, Telemarketer AI uniquely had higher acceptance for development over usage (; ). Additionally, confidence on using Telemarketer AI () was significantly () higher than that of Government Eligibility Interviewer AI (), which had the lowest confidence in usage. Near 0 mean for development of Lawyer AI () and usage for Telemarketer AI () suggest disagreement within judgments. Moreover, Elementary School Teacher AI uniquely unacceptable across both acceptability judgments (, ) underscoring a unique characteristic perhaps related to care. See Figure 3 for further details
Personal Use Cases
In personal use scenarios, Digital Medical Advice AI (; ), reflecting high risk level, consistently had lower acceptance across judgment types, compared to all other use cases. However, Customized Lifestyle Coach AI had the lowest confidence across both judgments (; ). Nutrition Optimizer (; ) had the highest mean acceptance across the two acceptability judgments. Interestingly, unlike the professional use cases, which had slightly higher acceptance for usage, personal use cases had lower acceptance for usage in general compared to development. Notably, many personal use cases had substantially high acceptance, especially compared to professional use cases.
| DEV() | USAGE() | |||
| Use Case Variation | Judgment | Confidence | Judgment | Confidence |
| Professional | ||||
| Education Level | ||||
| Intercept | ||||
| Personal | ||||
| EU AI Risk Level | ||||
| Intercept | ||||
5.1.2. Use Case Variations
When selecting use cases, we used two underlying variations: entry level of education required for professional use cases and EU AI risk levels for personal use cases. As risk levels and required education increased, we observe consistent negative effects on judgments, with personal use cases (; ) showing stronger effects compared to professional scenarios (), where only development judgments were significantly associated. Again, acceptance for personal use cases (; ) were higher than professional use cases ( ) as conveyed by the intercepts. Confidence ratings remained consistently high across all conditions (intercepts ¿ 3.97), though they showed a small but significant decrease with increasing risk levels in personal use cases (; ), while professional use cases showed no significant impact on confidence. Thus, while education level required for entry does show some significant effect, the EU AI risk level had consistent effects with higher significance.
5.1.3. Disagreements
Observing the standard deviation and visualization of the distribution shows further understanding of possible disagreements among use cases. We compare the judgments weighted by confidence to understand not only the differences in judgment but also their strength. Interestingly, the use cases with four highest disagreements in both judgments were all professional uses in order of Telemarketer (; ), Elementary School Teacher (; ), Lawyer (; ), and Government Eligibility Interviewer AI (; ). These four use cases were followed by Digital Medical Advice AI (, ). The use cases with the lowest disagreements were surprisingly Nutrition Optimizer (; ) followed by IT Support Specialist AI (; ). These results underscore the general controversy of professional usages but also shows that acceptability is highly use case dependent.
5.2. RQ2. Impact of Personal Factors on Acceptability Judgment
5.2.1. Demographic Factors
| DEV ( (SE)) | USAGE ( (SE)) | |||||
| Demographics | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
| Age | ||||||
| 25-34 | ||||||
| 55-64 | ||||||
| 25-34 | ||||||
| Race | ||||||
| Asian | ||||||
| Black | ||||||
| Mixed | ||||||
| Asian | ||||||
| Asian | ||||||
| Black | ||||||
| Gender | ||||||
| Non-male | ||||||
| Non-male | ||||||
| Political View | ||||||
| Str. liberal | ||||||
| Str. Liberal | ||||||
| Str. Liberal | ||||||
| Liberal | ||||||
| Education | ||||||
| Advanced | ||||||
| Employment | ||||||
| 40+ hrs | ||||||
| Discrimination | ||||||
| High | ||||||
In analyzing the demographic factors influencing judgments on the development and usage of AI use cases, several significant trends emerged from the data. First, we observed that the demographic variables had less than 0.5 correlation, except for age 65+ and Retired employment status. Across both categories of use cases, age was positively associated with confidence in acceptability judgment of usage for age 25-34 (, ) and 55-64 (, ). Race also had notable influences; specifically, Asian participants exhibited significantly lower confidence in both development and usage judgments (, ; , ), particularly in professional contexts.
Gender emerged as a crucial determinant, with non-male participants consistently showing negative judgments across both development (, ) and usage (, ), indicating potential discrepancies in perception or experience with AI applications. Liberal views, especially among those identifying as strongly liberal, were associated with negative judgments across both categories of use cases (, ; , ), suggesting a skeptical stance towards AI’s prevalence and role. Employment hours also contributed, with individuals working 40+ hours per week displaying a positive association with development judgments (, ), suggesting more exposure or reliance on AI use cases. High experience of discrimination chronicity was significantly related to lower acceptance of development, , . These findings highlight the significant role of demographic factors in shaping perceptions and attitudes toward AI technologies. See Table 23 in the Appendix for ANOVA results.
| DEV ( (SE)) | USAGE ( (SE)) | |||||
| AI Literacy | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
| Professional | ||||||
| AI Ethics | ||||||
| Gen AI Usage Freq. | ||||||
| Gen AI Limit. Familiarity | ||||||
| Personal | ||||||
| AI Skills | ||||||
| AI Ethics | ||||||
| Gen AI Usage Freq. | ||||||
| Gen AI Limit. Familiarity | ||||||
5.2.2. AI Literacy
We identified a correlation greater than 0.5 among three AI literacy aspects: awareness, usage, and evaluation. These were aggregated into a single factor, AI Skills. As shown in Table 4, understanding of AI Ethics was associated with lower acceptability for both personal (, ; , ) and professional (). However, across both categories of use cases, high Generative AI Usage Frequency resulted in higher acceptance for both professional (, ; , ) and usage acceptance (, ; , ). Notably, for personal use cases, AI Skills was positively associated with confidence of judgments (, for both development and usage) and judgment weighted by confidence (, ), while Generative AI Limitation Familiarity was positively associated with confidence for professional usage (, ; , for usage). These results implicate that different understandings of and experiences with AI can impact judgments of acceptability, corroborating previous findings (Kramer et al., 2018).
5.3. RQ3. Factors in Participant Rationale
To provide deeper insights to our analyses of acceptability judgments, we also examined open-text rationales (Q3) for judgments of development (Q1) and conditions for switching their decisions (Q4).
5.3.1. Decision-making Types
As we defined in § 4.2, we focus on two distinct reasoning types for decision-making, distinguished by consideration of outcome versus consideration of value inherent in action: cost-benefit reasoning (e.g., “it gives more people access to medical advice and treatment”, P365) and rule-based reasoning (e.g., should not be developed because “human interaction is better”, P249). As shown in Figure 5, generally, participants used more cost-benefit reasoning with highest percentage for IT Support Specialist (91.0%) and Nutrition Optimizer (92.8%). This result is particularly interesting as we observed these two use cases to have the lowest disagreement (see § 5.1.1) and signifies that perhaps unified reasoning type leads to less disagreement.
On the other hand, rationales for Elementary School Teacher AI contained most percentage of rule-based reasoning (30.0%) followed by Lawyer AI (22.0%). Interestingly, Elementary School Teacher AI and Lawyer AI were the use cases with the lowest development acceptability. These results suggest that less acceptable AI use cases might trigger more rule-based reasoning, which could, however, be due to the lack of consensus and rules around AI, especially those that are positive. this suggests the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable. These results suggest that the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable.
5.3.2. Moral Foundations
Beyond reasoning types, we explored moral foundations to provide insights into what values are relevant for AI use case decisions (Figure 6). For example, P12 responded that Elementary School Teacher AI use case should be developed because it “could give elementary schooling to children who are bed ridden…”, which was annotated with both values of Care (focusing on the well-being and nurturing of bed-ridden children) and Fairness (focusing on fair access to education). Upon analysis, we observe that Care (i.e., dislike of pain of others, feelings of empathy and compassion toward others) was the most prevalent moral value in participants’ rationales across the use cases. Of note, Care could impact the acceptability of AI use cases in both directions, as conveyed by P385 who noted that Customized Lifestyle Coach AI should be developed because “it may help improve some people’s health” but would change their decision if “it caused harm to even one person.” While our choice of medical domain in personal usage could have impacted the distribution, care was still the most prevalent considering only professional use cases.
Participants’ rationales evoked fairness value more than care for two use cases, Lawyer AI (0.62) and Government Eligibility Interviewer (0.74) respectively. These results could be due to the characteristics of the use cases, such as their main purpose and function, as noted by P88, Government Eligibility Interviewer should be developed because “it might be less biased and therefore more fair in it decisions (sic)”. Authority was most apparent in participant rationales for Lawyer AI (0.37) and Purity for Flavorful Swaps (0.35).
5.3.3. Switching Conditions
We have thus far explored firmness of the decision through levels of confidence provided by users. We further explore the flexibility of participants’ judgments to understand possible mitigation of disagreements through concerns expressed in conditions under which they would switch their decisions (Figure 7). Overall, Functionality (53.3%; e.g., Medical Adivce AI should not be developed if it “consistently or had a high percentage of failure to diagnose correctly.”, P373) was the most commonly noted concern that would switch participants decision. This was followed by Societal Impact (42.3%; e.g., Lawyer AI should not be developed if “it puts too many human lawyers out of work”, P97), Usage (31.8%; e.g., Government Eligibility Interviewer should be developed if “it was only used to read and screen applications but not for making decisions”), Not Applicable (0.02%; e.g., will not change decision).
Interestingly, Societal Impact (50.6%) was more prevalent followed by Functionality (46.4%) for professional use cases whereas Functionality (54.4%) was the most frequently mentioned concern for personal use cases. One of the reasons for frequent mention of Societal Impact in professional use cases could be due to concerns of labor replacement for professional use cases: as described by P296, if Elementary School Teacher AI “was to replace teachers with the ai to save money”, they would switch decision from should be developed to should not be developed. Moreover, we observe that Societal Impact was the most prevalent concern for those (31%) who thought the use case was unacceptable across all use cases and was especially prevalent for Elementary School Teacher AI (0.47). These diverging results show the importance of understanding granular concerns by use cases to effectively address the relevant issues. Participants’ responses that were identified as being both rule-based and concerned about societal impact expressed rules such as necessity of humanness as expressed by P22, “a human touch is 100% necessary” when discussing Elementary School AI, whether it be due to belief that “AI would lack empathy” (P5) or that “humans need human interactions” (P44).
| DEV ( (SE)) | USAGE ( (SE)) | |||||
| Rationale Factors | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
| Reasoning Type | ||||||
| Cost-benefit | ||||||
| Rule-based | ||||||
| Moral Value | ||||||
| Care | ||||||
| Fairness | ||||||
| Purity | ||||||
| Authority | ||||||
| Switching Condition | ||||||
| Functionality | ||||||
| Usage | ||||||
| Societal Impact | ||||||
5.3.4. Rationale Factors and Judgment
The results of analyzing the influence of rationale factors on judgments of acceptability indicated several significant effects (Table 5). We first verified that the factors had less than 0.5 correlation with each other. For acceptability of development, use of Cost-benefit reasoning in rationale was associated with increased acceptability (), while Rule-based reasoning was negatively associated with acceptability (). Interestingly, for usage, Cost-benefit reasoning was negatively associated with acceptability () but positively for judgment weighted by confidence (). Rule-based reasoning indicated lower usage acceptability for judgment weighted by confidence, consistent with acceptability of development. In the moral value category, Fairness in rationales had positive associations with acceptance (, ; judgment weighted by confidence).
Analyzing switching conditions showed that participants who answered negatively to the development of use cases mentioned the concerns about Functionality () and Societal Impact to switch their decisions (). However, those who were positive towards development of use case indicated emphasis on Usage () as a condition to reverse their decisions. Thus, our findings highlight diverse perspectives on requirements and concerns of AI use cases, especially with varying perception of acceptability.
6. Conclusion and Discussion
We conducted a study to understand how and why laypeople perceive various AI use cases as acceptable or not. To achieve this, we developed a survey that gathered judgments and reasoning processes from 197 participants who were demographically diverse and had varying levels of experience with AI. Participants were asked to provide their judgments on the acceptability of AI use cases, along with rationales for their decisions (e.g., ”Should / Should not be developed, because…”) and conditions that might change their decisions (e.g., ”I would switch my decision if…”). The survey covered ten different AI use cases, spanning both personal and professional domains, and included varying levels of risk. Our findings revealed significant variation in the acceptability judgments and reasoning factors based on the domain, risk level, and participants’ attributes, such as AI literacy and gender. We discuss the implications of these findings below.
Use Case Perceptions and Disagreements
In our study, we explored the varying acceptability of AI across different use cases. Generally, acceptance was lower in scenarios with higher educational requirements and greater EU AI risk levels. Professional use cases displayed more variability, notably with Elementary School Teacher AI, which was uniquely unacceptable. This underscores the necessity for further research into how AI should be developed and integrated, as well as what skills it should have, particularly in fields where empathy and care are crucial (Wu et al., 2024; Kawakami et al., 2024; Borg and Read, 2024). In addition, prior FAccT research have also highlighted how AI practitioners desire understanding lay people’s perception on AI fairness in specific use cases (Sonboli et al., 2021; Deng et al., 2022; Smith et al., 2023; Deng et al., 2023). Drawing from prior HCI and AI research (Deng et al., 2025; Lee et al., 2019; Cheng et al., 2019), future FAccT researchers and practitioners should explore how to meaningfully connect lay people’s use case perceptions with AI developers’ workflows.
While prior research has emphasized understanding AI consequences (Kieslich et al., 2024) and providing tools and processes to uncover impact (Wang et al., 2024; Buçinca et al., 2023; Deng et al., 2024), our findings reveal a greater presence of rule-based reasoning in contentious use cases, suggesting a need for diverse approaches to understanding AI beyond mere consequence anticipation. Moreover, while care was generally predominant, we observed that fairness gained prominence in Lawyer AI and Government Eligibility Interviewer AI. This variability underscores the importance of considering values in AI evaluation and training (Barocas et al., 2021; Bhardwaj et al., 2024), rather than solely emphasizing functionality, which is the current trend in AI research (Birhane et al., 2022). Additionally, societal impact considerations were more evident in unacceptable use cases, emphasizing the necessity for implementing safety guardrails when deploying AI with significant social implications (Solaiman, 2023).
Demographics and AI Literacy
In line with prior work (Kingsley et al., 2024; Mun et al., 2024), our results highlighted significant differences among demographic groups and perceived acceptance of use cases, especially for professional use (§5.2). Non-majority demographic groups, especially non-male gender groups, found both personal and professional use cases less acceptable. Those experiencing high discrimination chronicity also found professional use less acceptable. Our findings provide future FAccT research with valuable empirical insights on AI integration in contexts such as workplace, where marginalized worker’s agency, earning, and occupational well-being are disproportionately affected (Ming et al., 2024; Alcover et al., 2021).
Furthermore, our work highlighted a potential polarization on perceptions of AI among workers as those with 40+ hours employment and those who had advanced degrees were more positive towards AI use cases, suggesting that the relationship stakeholders have to AI and jobs might influence acceptability. Such a concern was expressed by one of our participants who opposed development of Telemarketer AI because it “overlaps with my industry, and hence serves as a threat to my job security” (P35). Thus, our results corroborate the need to further explore methods to include diverse workers and various stakeholders into the discussion of workplace AI integration and development (Fox et al., 2020; Cheon, 2023). We also found that frequent AI usage increased acceptance, while understanding AI ethics and limitations decreased acceptance. This suggests that balanced AI awareness and education, encompassing usage, skills, and ethics, could guide and improve decision-making (Raji et al., 2021), e.g., through educational interventions targeting AI skills and ethical implication literacy (e.g., (Wong and Nguyen, 2021; Shen et al., 2021)).
Rationales
Through analyzing participants’ rationales, we observed an interesting pattern with higher cost-benefit reasoning use cases with least disagreement and more prevalence of rule-based reasoning for those with higher disagreement (§5.3). These results suggest different valuation systems employed by participants perhaps leading to different conclusions and suggests some use cases have beyond simple utilitarian implication for society, which should be explored more carefully in future studies. Building upon our empirical findings, future work could develop tools and interventions which encourage specific types of acceptability reasoning such as rule and value based (Sorensen et al., 2024) or cost-benefit analyses (Li et al., 2024).
However, as our study was limited to the two reasoning type categories, expanding this analysis would be essential for future work including finding ways to classify what features people are considering in their decisions, how the weights on those features impact what kind of decision-making strategy they will use, and whether there are other ways to understand their decision strategies beyond our current classification. Future FAccT research can build upon these further understandings to guide policy making and consensus building. For example, in addition to conducting surveys with single participant, future work can explore how group discussions and deliberations shape communities’ collective understanding of AI impacts (e.g., (Kuo et al., 2024; Lee et al., 2019; DeVos et al., 2022; Gordon et al., 2022; Zhang et al., 2023))
Appendix A Additional Survey Details
A.1. Survey Questions
In our survey we ask participants to read one or more descriptions of AI use cases and to make two judgments: 1) “Do you think a technology like this should exist?” (Q1) and 2) “If the ¡use case¿ exists, would you use its services?” (Q5). A question to indicate their level of confidence is asked following each question (Q5, Q6). The participants are asked to both elaborate on their decisions (Q3) and specify the conditions under which they would switch their decisions (Q4). The detailed wordings for the questions are shown in Table 6.
Following the main use case questions in both the main and second study, we also asked participants questions about their demographics and literacy levels in AI, and the questions can be found in Table 7 and 8 respectively.
Lastly, while not included in the main text, we asked participants 3 questionnaires about decision making styles to explore the relationship between several decision making styles and the actual decisions of the participants. These included: (1) Oxford Utilitarianism Scale, (2) Toronto Empathy Questionnaire and (3) Moral Foundations Questionnaire. The decision making style questions can be found in Table 9, 10 and 11 respectively.
| Question ID | Question | Answer Type |
| AI Perception Question (Before) | ||
| AI Perception Before | Overall, how does the growing presence of artificial intelligence (AI) in daily life and society make you feel? | 5 Point Likert Scale |
| Part 1 Specific Questions | ||
| UCX-1 | Do you think a technology like this should be developed? | Yes/No |
| UCX-2 | How confident are you in your above answer? | 5 Point Likert Scale |
| UCX-3Y | Please complete the following: [Use Case] should be developed because… | Text |
| UCX-3N | Please complete the following: [Use Case] should not be developed because… | Text |
| UCX-4Y | Under what circumstances would you switch your decision from [UCX-2 Answer] should be developed to should not be developed? | Text |
| UCX-4N | Under what circumstances would you switch your decision from [UCX-2 Answer] should not be developed to should be developed? | Text |
| UCX-5 | If [Use Case] exists, would you ever use its services (answer yes, even if you think you would use it very infrequently)? | Yes/No |
| UCX-6 | How confident are you in your above answer? | 5 Point Likert Scale |
| AI Perception Question (After) | ||
| AI Perception After | Before we continue, we’d like to get your thoughts on AI one more time. Overall, how does the growing presence of artificial intelligence (AI) in daily life and society make you feel? | 5 Point Likert Scale |
| Question ID | Question |
| D-Q1 | How old are you? |
| D-Q2 | Choose one or more races that you consider yourself to be |
| D-Q3 | Do you identify as transgender? |
| D-Q4 | How would you describe your gender identity? |
| D-Q5 | How would you describe your sexual orientation? |
| D-Q6 | What is your present religion or religiosity, if any? |
| D-Q7 | In general, would you describe your political views as… |
| D-Q8 | What is the highest level of education you have completed? |
| D-Q9 | In which country have you lived in the longest? |
| D-Q10 | What other countries have you lived in for at least 6 months? |
| D-Q11 | Which of the following categories best describe your employment status? |
| D-Q12 | How would you describe the industry your job would be in? (Select all that apply) |
| D-Q13 | Do you identify with any minority, disadvantaged, demographic, or other specific groups? If so, which one(s)? (E.g., racial, gender identity, sexuality, disability, immigrant, veteran, etc.); use commas to separate groups. |
| D-Q14 | (Optional) What are some things that you are most concerned about lately? |
| In your day-to-day life how often have any of the following things happened to you? You are treated with less courtesy or respect than other people | |
| In your day-to-day life how often have any of the following things happened to you? You receive poorer service than other people at restaurants or stores | |
| In your day-to-day life how often have any of the following things happened to you? People act as if they think you are not smart | |
| In your day-to-day life how often have any of the following things happened to you? People act as if they are afraid of you | |
| In your day-to-day life how often have any of the following things happened to you? You are threatened or harassed | |
| D-Q16 | If the answer to Q15 is ”A few times a year” or more frequently to at least one of the statements, what do you think is the main reason for these experiences? (Select all that apply) |
| Question ID | Question |
| AI-Q1 | I can identify the AI technology employed in the applications and products I use. |
| AI-Q2 | I can skillfully use AI applications or products to help me with my daily work. |
| AI-Q3 | I can choose the most appropriate AI application or product from a variety for a particular task. |
| AI-Q4 | I always comply with ethical principles when using AI applications or products. |
| AI-Q5 | I am never alert to privacy and information security issues when using AI applications or products. |
| AI-Q6 | I am always alert to the abuse of AI technology. |
| AI-Q7 | How frequently do you use generative AI (i.e., artificial intelligence that is capable of producing high quality texts, images, etc. in response to prompts) products such as ChatGPT, Bard, DALL·E 2, Claude, etc.? |
| AI-Q8 | How familiar are you with limitations and shortcomings of generative AI? |
| Question ID | Question |
| Util1 | If the only way to save another person’s life during an emergency is to sacrifice one’s own leg, then one is morally required to make this sacrifice. |
| Util2 | It is morally right to harm an innocent person if harming them is a necessary means to helping several other innocent people. |
| Util3 | From a moral point of view, we should feel obliged to give one of our kidneys to a person with kidney failure since we don’t need two kidneys to survive, but really only one to be healthy. |
| Util4 | If the only way to ensure the overall well-being and happiness of the people is through the use of political oppression for a short, limited period, then political oppression should be used. |
| Util5 | From a moral perspective, people should care about the well-being of all human beings on the planet equally; they should not favor the well-being of people who are especially close to them either physically or emotionally |
| Util6 | It is permissible to torture an innocent person if this would be necessary to provide information to prevent a bomb going off that would kill hundreds of people. |
| Util7 | It is just as wrong to fail to help someone as it is to actively harm them yourself. |
| Util8 | Sometimes it is morally necessary for innocent people to die as collateral damage if more people are saved overall. |
| Util9 | It is morally wrong to keep money that one doesn’t really need if one can donate it to causes that provide effective help to those who will benefit a great deal. |
| Question ID | Question |
| Empathy1 | When someone else is feeling excited, I tend to get excited too. |
| Empathy2 | Other people’s misfortunes do not disturb me a great deal. |
| Empathy3 | It upsets me to see someone being treated disrespectfully. |
| Empathy4 | I remain unaffected when someone close to me is happy. |
| Empathy5 | I enjoy making other people feel better. |
| Empathy6 | I have tender, concerned feelings for people less fortunate than me. |
| Empathy7 | When a friend starts to talk about his/her problems, I try to steer the conversation towards something else. |
| Empathy8 | I can tell when others are sad even when they do not say anything. |
| Empathy9 | I find that I am “in tune” with other people’s moods. |
| Empathy10 | I do not feel sympathy for people who cause their own serious illnesses. |
| Empathy11 | I become irritated when someone cries. |
| Empathy12 | I am not really interested in how other people feel. |
| Empathy13 | I get a strong urge to help when I see someone who is upset. |
| Empathy14 | When I see someone being treated unfairly, I do not feel very much pity for them. |
| Empathy15 | I find it silly for people to cry out of happiness. |
| Empathy16 | When I see someone being taken advantage of, I feel kind of protective towards him/her. |
| Question ID | Question |
Moral Foundation Questionnaire (First Half)
When you decide whether something is right or wrong, to what extent is the following consideration relevant to your thinking? |
| MFQ 1 | Whether or not someone suffered emotionally |
| MFQ 2 | Whether or not some people were treated differently than others |
| MFQ 3 | Whether or not someone’s action showed love for his or her country |
| MFQ 4 | Whether or not someone showed a lack of respect for authority |
| MFQ 5 | Whether or not someone violated standards of purity and decency |
| MFQ 6 | Whether or not someone was good at math |
| MFQ 7 | Whether or not someone cared for someone weak or vulnerable |
| MFQ 8 | Whether or not someone acted unfairly |
| MFQ 9 | Whether or not someone did something to betray his or her group |
| MFQ 10 | Whether or not someone conformed to the traditions of society |
| MFQ 11 | Whether or not someone did something disgusting |
| Moral Foundation Questionnaire (Second Half) | |
| MFQ 12 | Compassion for those who are suffering is the most crucial virtue. |
| MFQ 13 | When the government makes laws, the number one principle should be ensuring that everyone is treated fairly. |
| MFQ 14 | I am proud of my country’s history. |
| MFQ 15 | Respect for authority is something all children need to learn. |
| MFQ 16 | People should not do things that are disgusting, even if no one is harmed. |
| MFQ 17 | It is better to do good than to do bad. |
| MFQ 18 | One of the worst things a person could do is hurt a defenseless animal. |
| MFQ 19 | Justice is the most important requirement for a society. |
| MFQ 20 | People should be loyal to their family members, even when they have done something wrong. |
| MFQ 21 | Men and women each have different roles to play in society. |
| MFQ 22 | I would call some acts wrong on the grounds that they are unnatural. |
A.2. Participant Details
The demographics of the participants for our study is shown in Tables 13 to 35. There was a fairly balanced distribution of participants across the different age groups, although there was a slightly higher proportion of participants in the 25-34 years old and 45-54 years old age ranges. In terms of racial distribution, there were more White/Caucasian participants compared to the other races. The gender distribution was relatively balanced in terms of males vs non-males. The participants were mostly employed or looking for work and a majority of them also had at least some form of college education. Most participants identified as liberal in terms of political leaning. Participants’ AI literacy scores are shown in Table 19 and AI Ethics score are shown in Table 20.
Participants were allocated 5 use cases from one of the scenarios and the allocation between the 2 scenarios are well-balanced and can be found in Table 12.
| Use Case | Participants Allocated |
| Personal Use Cases | |
| Digital Medical Advice | 97 |
| Customized Lifestyle Coach | |
| Personal Health Research | |
| Nutrition Optimizer | |
| Flavorful Swaps | |
| Labor Replacement Use Cases | |
| Lawyer | 100 |
| Elementary School Teacher | |
| IT Support Specialist | |
| Government Eligibility Interviewer | |
| Telemarketer |
| Racial Identity | (N) (%) | Age | N (%) | Gender Identity | N (%) | Education | N (%) |
| White or Caucasian | 33 (33.0) | 18-24 | 11 (11.0) | Man | 49 (49.0) | Bachelor’s degree | 36 (36.0) |
| Black or African American | 23 (23.0) | 45-54 | 27 (27.0) | Non-male | 51 (51.0) | Graduate degree∗ | 18 (18.0) |
| Asian | 21 (21.0) | 25-34 | 22 (22.0) | Some college ∗ | 17 (17.0) | ||
| Mixed | 13 (13.0) | 55-64 | 20 (20.0) | High school diploma∗ | 16 (16.0) | ||
| Other | 10 (10.0) | 35-44 | 14 (14.0) | Associates degree∗ | 13 (13.0) | ||
| 2 (0.7) | 65+ | 6 (6.0) | Some high school∗ | 0 (0.0) |
| Minority/Disadvantaged Group | (N) (%) | Transgender | N (%) | Sexuality | N (%) | Political Leaning | N (%) |
| No | 68 (68.0) | No | 97 (97.0) | Heterosexual | 78 (78.0) | Liberal | 34 (34.0) |
| Yes | 32 (32.0) | Yes | 2 (2.0) | Others | 22 (22.0) | Moderate | 23 (23.0) |
| Prefer not to say | 1 (1.0) | Strongly liberal | 20 (20.0) | ||||
| Conservative | 18 (18.0) | ||||||
| Strongly conservative | 4 (4.0) | ||||||
| Prefer not to say | 1 (1.0) |
| Longest Residence | (N) (%) | Employment | N (%) | Occupation (Top 10) | N (%) | Religion | N (%) |
| United States of America | 97 (97.0) | Employed, 40+ | 53 (53.0) | Other | 35 (35.0) | Christian | 29 (29.0) |
| Others | 3 (3.0) | Employed, 1-39 | 16 (16.0) | Prefer not to answer | 10 (10.0) | Agnostic | 20 (20.0) |
| Retired | 9 (9.0) | Health Care and Social Assistance | 10 (10.0) | Atheist | 15 (15.0) | ||
| Not employed, looking for work | 7 (7.0) | Information | 10 (10.0) | Nothing in particular | 13 (13.0) | ||
| Disabled, not able to work | 5 (5.0) | Manufacturing | 7 (7.0) | Catholic | 11 (11.0) | ||
| Not employed, NOT looking for work | 4 (4.0) | Professional, Scientific, and Technical Services | 7 (7.0) | Muslim | 5 (5.0) | ||
| Other: please specify | 4 (4.0) | Arts, Entertainment, and Recreation | 6 (6.0) | Hindu | 3 (3.0) | ||
| Prefer not to disclose | 2 (2.0) | Retail Trade | 6 (6.0) | Something else, Specify | 2 (2.0) | ||
| Finance and Insurance | 5 (5.0) | Jewish | 1 (1.0) | ||||
| Transportation and Warehousing, and Utilities | 4 (4.0) | Buddhist | 1 (1.0) |
| Racial Identity | (N) (%) | Age | N (%) | Gender Identity | N (%) | Education | N (%) |
| White or Caucasian | 29 (29.9) | 45-54 | 30 (30.9) | Man | 50 (51.5) | Bachelor’s degree | 40 (41.2) |
| Black or African American | 26 (26.8) | 25-34 | 26 (26.5) | Non-male | 47 (48.5) | Some college ∗ | 21 (21.6) |
| Asian | 20 (20.6) | 55-64 | 14 (14.4) | Graduate degree∗ | 14 (14.4) | ||
| Other | 14 (14.4) | 35-44 | 13 (13.4) | High school diploma∗ | 13 (13.4) | ||
| Mixed | 8 (8.2) | 18-24 | 9 (9.3) | Associates degree∗ | 8 (8.2) | ||
| 2 (0.7) | 65+ | 4 (4.1) | Some high school∗ | 1 (1.0) | |||
| Prefer not to disclose | 1 (1.0) |
| Minority/Disadvantaged Group | (N) (%) | Transgender | N (%) | Sexuality | N (%) | Political Leaning | N (%) |
| No | 51 (52.6) | No | 94 (96.9) | Heterosexual | 75 (77.3) | Liberal | 34 (35.1) |
| Yes | 46 (47.4) | Yes | 3 (3.1) | Others | 22 (22.7) | Moderate | 31 (32.0) |
| Prefer not to say | 0 (0.0) | Strongly liberal | 12 (12.4) | ||||
| Conservative | 10 (10.3) | ||||||
| Strongly conservative | 9 (9.3) | ||||||
| Prefer not to say | 1 (1.0) |
| Longest Residence | (N) (%) | Employment | N (%) | Occupation (Top 10) | N (%) | Religion | N (%) |
| United States of America | 93 (95.9) | Employed, 40+ | 46 (47.4) | Other | 36 (35.6) | Christian | 40 (40.8) |
| Others | 4 (4.1) | Employed, 1-39 | 22 (22.7) | Health Care and Social Assistance | 11 (11.3) | Catholic | 16 (16.3) |
| Not employed, looking for work | 13 (13.4) | Prefer not to answer | 10 (10.3) | Agnostic | 15 (15.3) | ||
| Not employed, NOT looking for work | 4 (4.1) | Professional, Scientific, and Technical Services | 9 (9.3) | Nothing in particular | 11 (11.2) | ||
| Disabled, not able to work | 4 (4.1) | Educational Services | 9 (9.3) | Atheist | 5 (5.1) | ||
| Other: please specify | 4 (4.1) | Finance and Insurance | 8 (8.2) | Something else, Specify | 5 (5.1) | ||
| Retired | 3 (3.1) | Arts, Entertainment, and Recreation | 5 (5.2) | Buddhist | 3 (3.1) | ||
| Prefer not to disclose | 1 (1.0) | Manufacturing | 5 (5.2) | Muslim | 1 (1.0) | ||
| Retail Trade | 4 (4.1) | Jewish | 1 (1.0) | ||||
| Accommodation and Food Services | 4 (4.1) | Hindu | 1 (1.0) |
| Score | AI Awareness | AI Usage | AI Evaluation | Gen AI Usage Freq. | Gen AI Limit. Familiarity |
| 1 | 25 | 15 | 30 | 35 | 55 |
| 2 | 40 | 60 | 75 | 200 | 320 |
| 3 | 75 | 90 | 105 | 155 | 345 |
| 4 | 140 | 125 | 125 | 235 | 230 |
| 5 | 380 | 310 | 275 | 220 | 35 |
| 6 | 280 | 300 | 310 | 140 | — |
| 7 | 45 | 85 | 65 | — | — |
| Score | AI Ethics |
| 5 | 15 |
| 6 | 10 |
| 7 | 25 |
| 8 | 25 |
| 9 | 75 |
| 10 | 105 |
| 11 | 165 |
| 12 | 100 |
| 13 | 105 |
| 14 | 90 |
| 15 | 120 |
| 16 | 80 |
| 17 | 35 |
| 18 | 35 |
A.3. Open-text Annotation Dimensions
Reasoning Type
Inspired by previous works in moral psychology, we used two main reasoning types to characterize participants’ decision making pattern as expressed in their open-text answers: cost-benefit reasoning and rule-based reasoning (Cheung et al., 2024). These two reasoning types are rooted in two decision making processes in moral and wider decision making literature: utilitarian and deontological reasoning, respectively. Cost-benefit reasoning thus considers the possible outcomes and their expected utility or value when making decisions, and rule-based reasoning shows more inherent value in action or entities. See § 2.3 for further discussion.
Moral Values
To annotate which values were prevalent in participants’ considerations of use cases, we used five moral values: care, fairness, loyalty, authority, and purity (Graham et al., 2011, 2008). While these dimensions have been re-defined to include more diverse values from participants beyond WEIRD (white, educated, industrialized, rich, and democratic) (Atari et al., 2023), we used these five dimensions due to brevity of the questionnaire, which was used in the survey to provide importance of each values to participants.
Switching Conditions
We annotated concerns expressed in switching conditions using three categories: functionality (e.g., errors, bias in systems, limited capabilities), usage (e.g., errors, bias in systems, limited capabilities), and societal impact (e.g., job loss, over-reliance), inspired by harm taxonomy developed by Solaiman et al. and user concern annotation practice adopted by Mun et al..
Appendix B Open-text Annotation Details
B.1. Automatic Annotation
B.1.1. Methods
We used Open-AI’s o1-mini model with maximum tokens set to 1024 to control response length, use a temperature of 0.7 to manage randomness, and keep top_p at 1 with default settings for frequency and presence penalties at 0. Prompts will be released with data upon acceptance.
B.1.2. Results
Results for inter-rater reliability analysis of o1’s annotations are shown in Table 21.
| Category | AC1 | Interpretation | 95% CI | p-value | z | SE | PA | PE |
| Cost Benefit | 0.976 | Almost Perfect | (0.942, 1.000) | 0.0 | 56.9 | 0.0172 | 0.980 | 0.164 |
| Rule Based | 0.848 | Almost Perfect | (0.754, 0.943) | 0.0 | 17.8 | 0.0476 | 0.890 | 0.276 |
| Care | 0.427 | Moderate | (0.234, 0.619) | 4.40 | 0.0970 | 0.650 | 0.390 | |
| Fairness | 0.411 | Moderate | (0.224, 0.599) | 4.35 | 0.0945 | 0.690 | 0.474 | |
| Authority | 0.839 | Almost Perfect | (0.743, 0.935) | 0.0 | 17.3 | 0.0486 | 0.880 | 0.255 |
| Purity | 0.758 | Substantial | (0.639, 0.878) | 0.0 | 12.6 | 0.0603 | 0.820 | 0.255 |
| Functionality | 0.587 | Moderate | (0.425, 0.749) | 7.18 | 0.0818 | 0.790 | 0.492 | |
| Usage | 0.673 | Substantial | (0.525, 0.822) | 9.03 | 0.0746 | 0.820 | 0.449 | |
| Societal Impact | 0.595 | Moderate | (0.432, 0.759) | 7.23 | 0.0823 | 0.770 | 0.432 |
Appendix C Factors Impacting Acceptability Judgments
C.1. Use Case Factors
Additional analysis of use case factors showing distribution of judgments by use case sorted by standard deviation is shown in Figure 8. Table 22 shows analysis of use case effect using ANOVA.
| Acceptability | Aspect | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
| EXIST | Judgment | Category | 29.903 | 29.903 | 1 | 197 | 5.98e-11 *** |
| Use Case | 86.116 | 9.5684 | 9 | 641.38 | ¡ 2.2e-16 *** | ||
| Confidence | Category | 0.0017113 | 0.0017113 | 1 | 197 | 0.9563 | |
| Use Case | 13.519 | 1.5021 | 9 | 603.23 2.7243 | 0.004037 ** | ||
| USAGE | Judgment | Category | 8.4257 | 8.4257 | 1 | 197 | 0.0002488 *** |
| Use Case | 73.801 | 8.2001 | 9 | 610.34 | ¡ 2.2e-16 *** | ||
| Confidence | Category | 1.3444 | 1.3444 | 1 | 197 | 0.153 | |
| Use Case | 20.721 | 2.3023 | 9 | 603.36 | 0.0001783 *** |
C.2. Demographics Factors
Additional analysis using ANOVA for demographic factors is shown in Table 23.
| EXIST | USAGE | |||||
| Demographics | Judgment | Confidence | Judg.Conf. | Judgment | Confidence | JudgConf. |
| All | ||||||
| Gender | ||||||
| Race | ||||||
| Employment | ||||||
| Sexual Orientation | ||||||
| Professional | ||||||
| Race | ||||||
| Gender | ||||||
| Education | ||||||
| Discrimination | ||||||
| Personal | ||||||
| Race | ||||||
| Political View | ||||||
| Employment | ||||||
| EXIST ( (SE)) | USAGE ( (SE)) | |||||
| Decision Style Factors | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
| MFQ Care | ||||||
| MFQ Fairness | ||||||
| MFQ Loyalty | ||||||
| MFQ Authority | ||||||
| MFQ Purity | ||||||
| Empathy | ||||||
| InstrumentalHarm | ||||||
| ImpartialBenificence | ||||||
| AIC | ||||||
| BIC | ||||||
| Log Likelihood | ||||||
| Num. obs. | ||||||
| Num. groups: prolific_id | ||||||
| Num. groups: use_case | ||||||
| Var: prolific_id (Intercept) | ||||||
| Var: use_case (Intercept) | ||||||
| Var: Residual | ||||||
C.2.1. Questionnaires
Interestingly, only Loyalty had a significant effect on both existence () and usage () as shown in Table 24. Moreover, Empathy had a positive and marginally significant effect for usage (). However, Loyalty, as shown in Figure 6, does not appear as frequently in participants’ open text responses compared to values such as Care and Fairness and is the only value that did not have a significant association with use cases.
Appendix D Factors in Participant Rationale
D.1. Reasoning Types
We show the flow of participants’ decisions and corresponding rationales throughout use cases in Figure 9, which shows interesting distribution and switching of reasoning types, which would be interesting for future studies to consider. Moreover, Table 25 shows that there are almost no relation between reasoning types used by the participants and the decision-making style questionnaire results signifying that the reasoning types might be highly use-case specific rather than a character trait. It would be interesting to study the factors that actually influence the choice of reasoning types.
| Cost-benefit | Rule-based | |
| MFQ Care | ||
| MFQ Fairness | ||
| MFQ Loyalty | ||
| MFQ Authority | ||
| MFQ Purity | ||
| Empathy | ||
| InstrumentalHarm | ||
| ImpartialBenificence | ||
| AIC | ||
| BIC | ||
| Log Likelihood | ||
| Num. obs. | ||
| Num. groups: prolific_id | ||
| Num. groups: use_case | ||
| Var: prolific_id (Intercept) | ||
| Var: use_case (Intercept) | ||
| Var: Residual | ||
D.2. Impact of Rationale Factors on Judgment
We display the analysis results using ANOVA to understand the effect of rationale factors on judgment in Table 26.
| Acceptability | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
| EXIST | Judgment | |||||
| Cost Benefit | 2.941 | 2.941 | 1 | 955.32 | 0.001306 ** | |
| Rule Based | 7.187 | 7.187 | 1 | 941.74 | 5.562e-07 *** | |
| Fairness | 1.551 | 1.551 | 1 | 966.41 | 0.019414 * | |
| Authority | 1.044 | 1.044 | 1 | 961.50 | 0.055021 . | |
| Functionality | 1.137 | 1.137 | 1 | 964.09 | 0.045230 * | |
| Usage | 125.594 | 125.594 | 1 | 837.10 | ¡ 2.2e-16 *** | |
| Societal Impact | 1.873 | 1.873 | 1 | 967.89 | 0.010228 * | |
| Confidence | ||||||
| Rule Based | 5.7805 | 5.7805 | 1 | 856.38 | 0.001158 ** | |
| Care | 3.8723 | 3.8723 | 1 | 907.71 | 0.007762 ** | |
| Fairness | 5.7117 | 5.7117 | 1 | 899.21 | 0.001237 ** | |
| Authority | 2.2665 | 2.2665 | 1 | 851.91 | 0.041526 * | |
| Usage | 2.6254 | 2.6254 | 1 | 809.17 | 0.028303 * | |
| Judgment x Confidence | ||||||
| Cost Benefit | 48.04 | 48.04 | 1 | 940.91 | 0.0004985 *** | |
| Rule Based | 86.12 | 86.12 | 1 | 924.15 | 3.331e-06 *** | |
| Fairness | 31.87 | 31.87 | 1 | 972.26 | 0.0045236 ** | |
| Authority | 14.24 | 14.24 | 1 | 972.97 | 0.0574858 . | |
| Functionality | 18.99 | 18.99 | 1 | 971.66 | 0.0283005 * | |
| Usage | 2454.30 | 2454.30 | 1 | 888.43 | ¡ 2.2e-16 *** | |
| Societal Impact | 34.26 | 34.26 | 1 | 971.98 | 0.0032486 ** | |
| USAGE | Judgment | |||||
| Cost Benefit | 0.28 | 0.28 | 1 | 933.91 | 0.01742 * | |
| Usage | 491.35 | 491.35 | 1 | 943.16 | ¡ 2e-16 *** | |
| Confidence | ||||||
| Cost Benefit | 2.2380 | 2.2380 | 1 | 864.41 | 0.0555024 . | |
| Rule Based | 2.0230 | 2.0230 | 1 | 852.41 | 0.0686374 . | |
| Fairness | 4.2517 | 4.2517 | 1 | 895.25 | 0.0083622 ** | |
| Authority | 2.4525 | 2.4525 | 1 | 852.20 | 0.0450309 * | |
| Loyalty | 2.2751 | 2.2751 | 1 | 875.10 | 0.0535163 . | |
| Functionality | 2.6366 | 2.6366 | 1 | 897.34 | 0.0376892 * | |
| Usage | 9.2673 | 9.2673 | 1 | 817.52 | 0.0001033 *** | |
| Societal Impact | 2.1237 | 2.1237 | 1 | 913.41 | 0.0620930 . | |
| Judgment x Confidence | ||||||
| Cost Benefit | 106.358 | 106.358 | 1 | 874.26 | 8.333e-05 *** | |
| Rule Based | 33.512 | 33.512 | 1 | 859.58 | 0.026742 * | |
| Fairness | 89.120 | 89.120 | 1 | 930.00 | 0.000312 *** | |
| Societal Impact | 74.975 | 74.975 | 1 | 929.00 | 0.000938 *** | |
D.3. Factors Influencing Moral Foundations in Rationale
In Table 27 we display analysis result using linear mixed effects on factors that may influence moral foundations appealed to in participants’ rationales. We find greater relations with the use cases than personal factors.
| Care | Fairness | Purity | Loyalty | Authority | |
| (Intercept) (Telemarketer) | |||||
| MFQ_care | |||||
| MFQ_fairness | |||||
| MFQ_loyalty | |||||
| MFQ_authority | |||||
| MFQ_purity | |||||
| empathy_total | |||||
| InstrumentalHarm | |||||
| ImpartialBenificence | |||||
| Gov. Eligi. Interviewer | — | ||||
| IT Support Specialist | — | ||||
| Elementary School Teacher | |||||
| Lawyer | |||||
| Flavorful Swaps | |||||
| Nutrition Optimizer | — | ||||
| Personal Health Research | — | ||||
| Cust, Lifestyle Coach | |||||
| Digital Medical Advice | — | ||||
| AIC | |||||
| BIC | |||||
| Log Likelihood | |||||
| Num. obs. | |||||
| Num. groups: prolific_id | |||||
| Var: prolific_id (Intercept) | |||||
Appendix E Survey 2: Explicit Weighing of Harms and Benefits of Use Cases
Although not included in main text, we administered a variation of our main study where we asked participants to explicitly reason through harms and benefits. The decisions were measured before and after the explicit weighing of harms and benefits. However, we saw almost no effect.
E.1. Study Overview
To better understand the reasoning behind participants decisions about the judgment and usage of the use cases, we conducted a second study with 1 survey for each category (Labor Replacement Use Cases and Personal Use Cases). The second study includes an additional set of questions to elicit the harms and benefits of developing and not developing an use case to better understand the reasoning behind participants decisions. Furthermore, we asked participants the judgment and usage decision questions before and after the set of harms and benefits questions to see if listing out reasons about an use case would elicit any change in their decisions. The details of the second study can be found in E.2 and the results can be found in E.3.
E.2. Setup and Details
While these same set of questions are asked for all five use cases for our main study, in our second study, participants are randomly allocated a single use case. The second study differs from the main study with an initial set of judgment questions without open-text rationales (Q1 - Initial to Q4 - Initial), which are followed by explicit listing and weighing of the possible harms and benefits of the use case in the context of both developing and not developing the use case. We then again ask participants the same set of judgment questions along with the open-text questions to elaborate on their reasoning, similar to the main study. To understand how the judgment and usage decisions are affected by other factors, we asked the participants about their demographics, ai literacy levels and several other reasoning factors after the main set of questions, and these questions can be found in §A.1 The main questions for the second study can be found in Table 29. The participant demographics for the second study can be found in Tables 30 to 35. The distribution of each use case within each scenario (Labor Replacement Use Cases and Personal Use Cases) for the second study is relatively well-balanced and can be found in Table 28.
| Use Case | Participants Allocated |
| Personal Use Cases | |
| Digital Medical Advice | 20 |
| Customized Lifestyle Coach | 20 |
| Personal Health Research | 19 |
| Nutrition Optimizer | 21 |
| Flavorful Swaps | 17 |
| Labor Replacement Use Cases | |
| Lawyer | 20 |
| Elementary School Teacher | 22 |
| IT Support Specialist | 17 |
| Government Eligibility Interviewer | 19 |
| Telemarketer | 23 |
| Question ID | Question | Answer Type |
| AI Perception Question (Before) | ||
| AI Perception Before | Overall, how does the growing presence of artificial intelligence (AI) in daily life and society make you feel? | 5 Point Likert Scale |
| Initial Decision/Usage | ||
| Q1 - Initial | Do you think a technology like this should be developed? | Yes/No |
| Q2 - Initial | How confident are you in your above answer? | 5 Point Likert Scale |
| Q3 - Initial | If [Use Case] exists, would you ever use its services (answer yes, even if you think you would use it very infrequently)? | Yes/No |
| Q4 - Initial | How confident are you in your above answer? | 5 Point Likert Scale |
| Benefits of Developing Use Case | ||
| Q1 - BDev | How will [Use Case] positively impact individuals? | Text |
| Q2 - BDev | Which groups of people do you think would benefit the most from the above positive impacts? (You can list more than one group.) | Text |
| Q3 - BDev | How beneficial would [Use Case] be if it had the above positive impacts? | 9 Point Likert Scale |
| Malicious Uses of Developing Use Case | ||
| Q1 - HDev | Please complete the following: [Use Case] could have a negative impact if it was used to… | Text |
| Q1 - HDev | What would be the negative impact of the above malicious or unintended uses? | Text |
| Q2 - HDev | Which groups of people do you think would be harmed the most by the above malicious or unintended uses? (You can list more than one group.) | Text |
| Q3 - HDev | How harmful would [Use Case] be if it had the above negative impacts? | 9 Point Likert Scale |
| Failures of Developing Use Case | ||
| Q1 - HDevF | Please complete the following: If [Use Case] failed to do its intended task properly, fully, and accurately, it could have a negative impact if it… | Text |
| Q1 - HDevF | What would be the negative impact of those failure cases? | Text |
| Q2 - HDevF | Which groups of people do you think would be harmed the most by the above failure cases? (You can list more than one group.) | Text |
| Q3 - HDevF | How harmful would [Use Case] be if it had the above negative impacts? | 9 Point Likert Scale |
| Benefits of Not Developing Use Case | ||
| Q1 - BNonDev | Please complete the following: Not having [Use Case] would be beneficial because… | Text |
| Q2 - BNonDev | Which groups of people do you think would benefit the most by banning or not developing [Use Case]? (You can list more than one group.) | Text |
| Q3 - BNonDev | How beneficial would it be if [Use Case] was banned or not developed and it had the above positive impact? | 9 Point Likert Scale |
| Harms of Not Developing Use Case | ||
| Q1 - HNonDev | Please complete the following: Not having [Use Case] would be harmful because… | Text |
| Q2 - HNonDev | Which groups of people do you think would be harmed the most by banning or not developing [Use Case]? (You can list more than one group.) | Text |
| Q3 - HNonDev | How harmful would it be if [Use Case] was banned or not developed and it had the above negative impacts? | 9 Point Likert Scale |
| Final Decision/Usage | ||
| Q1 - Final | Do you think a technology like this should be developed? | Yes/No |
| Q2 - Final | How confident are you in your above answer? | 5 Point Likert Scale |
| Q3 - Final - Y | Please elaborate on your answer to the previous question: Do you think a technology like this should be developed?: [Q1 - Final Answer] | Text |
| Q3 - Final - N | Please elaborate on your answer to the previous question: Do you think a technology like this should be developed?: [Q1 - Final Answer] | Text |
| Q4 - Final - Y | Under what circumstances would you switch your decision from [Q1 - Final Answer] should be developed to should not be developed? | Text |
| Q4 - Final - N | Under what circumstances would you switch your decision from [Q1 - Final Answer] should not be developed to should be developed? | Text |
| Q5 - Final | If [Use Case] exists, would you ever use its services (answer yes, even if you think you would use it very infrequently)? | Yes/No |
| Q6 - Final | How confident are you in your above answer? | 5 Point Likert Scale |
| AI Perception Question (After) | ||
| AI Perception After | Before we continue, we’d like to get your thoughts on AI one more time. Overall, how does the growing presence of artificial intelligence (AI) in daily life and society make you feel? | 5 Point Likert Scale |
| Racial Identity | (N) (%) | Age | N (%) | Gender Identity | N (%) | Education | N (%) |
| White or Caucasian | 32 (31.4) | 45-54 | 32 (31.4) | Man | 49 (49.0) | Bachelor’s degree | 44 (43.1) |
| Black or African American | 25 (24.5) | 25-34 | 29 (28.4) | Non-male | 51 (51.0) | Graduate degree∗ | 18 (17.6) |
| Asian | 18 (17.6) | 35-44 | 12 (11.8) | Some college ∗ | 18 (17.6) | ||
| Mixed | 15 (14.7) | 55-64 | 12 (11.8) | High school diploma∗ | 15 (14.7) | ||
| Other | 12 (11.8) | 18-24 | 9 (8.8) | Associates degree∗ | 7 (6.9) | ||
| 2 (0.7) | 65+ | 8 (7.8) | Some high school∗ | 0 (0.0) |
| Minority/Disadvantaged Group | (N) (%) | Transgender | N (%) | Sexuality | N (%) | Political Leaning | N (%) |
| No | 56 (54.9) | No | 97 (95.1) | Heterosexual | 76 (74.5) | Liberal | 37 (36.3) |
| Yes | 46 (45.1) | Yes | 4 (3.9) | Others | 26 (25.5) | Moderate | 27 (26.5) |
| Prefer not to say | 1 (1.0) | Conservative | 17 (16.7) | ||||
| Strongly liberal | 16 (15.7) | ||||||
| Strongly conservative | 4 (3.9) | ||||||
| Prefer not to say | 1 (1.0) |
| Longest Residence | (N) (%) | Employment | N (%) | Occupation (Top 10) | N (%) | Religion | N (%) |
| United States of America | 96 (94.1) | Employed, 40+ | 44 (43.1) | Other | 34 (33.3) | Christian | 38 (37.3) |
| Others | 6 (5.9) | Employed, 1-39 | 28 (27.5) | Educational Services | 11 (10.8) | Agnostic | 19 (18.6) |
| Not employed, looking for work | 19 (18.6) | Health Care and Social Assistance | 10 (10.0) | Catholic | 18 (17.6) | ||
| Retired | 4 (3.9) | Information | 8 (7.8) | Nothing in particular | 12 (11.8) | ||
| Not employed, NOT looking for work | 3 (2.9) | Prefer not to answer | 8 (7.8) | Atheist | 7 (6.9) | ||
| Other: please specify | 3 (2.9) | Retail Trade | 7 (6.9) | Muslim | 3 (2.9) | ||
| Disabled, not able to work | 1 (1.0) | Finance and Insurance | 7 (6.9) | Something else, Specify | 3 (2.9) | ||
| Prefer not to disclose | 0 (0.0) | Professional, Scientific, and Technical Services | 6 (5.9) | Jewish | 1 (1.0) | ||
| Manufacturing | 6 (5.9) | Hindu | 1 (1.0) | ||||
| Administrative and support and waste management services | 5 (4.9) | Buddhist | 0 (0.0) |
| Racial Identity | (N) (%) | Age | N (%) | Gender Identity | N (%) | Education | N (%) |
| White or Caucasian | 35 (36.1) | 45-54 | 29 (29.9) | Non-male | 50 (51.5) | Bachelor’s degree | 33 (34.0) |
| Black or African American | 22 (22.7) | 25-34 | 22 (22.7) | Man | 47 (48.5) | Graduate degree∗ | 24 (24.7) |
| Asian | 19 (19.6) | 55-64 | 17 (17.5) | Some college ∗ | 18 (18.6) | ||
| Mixed | 13 (13.4) | 35-44 | 17 (17.5) | High school diploma∗ | 12 (12.4) | ||
| Other | 8 (8.2) | 18-24 | 6 (6.2) | Associates degree∗ | 10 (10.3) | ||
| 2 (0.7) | 65+ | 6 (6.2) | Some high school∗ | 0 (0.0) | |||
| Prefer not to disclose | 0 (0.0) |
| Minority/Disadvantaged Group | (N) (%) | Transgender | N (%) | Sexuality | N (%) | Political Leaning | N (%) |
| No | 50 (51.5) | No | 93 (95.9) | Heterosexual | 76 (78.4) | Liberal | 34 (35.1) |
| Yes | 47 (48.5) | Yes | 4 (4.1) | Others | 21 (21.6) | Moderate | 26 (26.8) |
| Prefer not to say | 0 (0.0) | Strongly liberal | 17 (17.5) | ||||
| Conservative | 13 (13.4) | ||||||
| Strongly conservative | 6 (6.2) | ||||||
| Prefer not to say | 1 (1.0) |
| Longest Residence | (N) (%) | Employment | N (%) | Occupation (Top 10) | N (%) | Religion | N (%) |
| United States of America | 95 (97.9) | Employed, 40+ | 44 (45.4) | Other | 36 (37.1) | Christian | 43 (44.3) |
| Others | 2 (2.1) | Employed, 1-39 | 23 (23.7) | Health Care and Social Assistance | 13 (13.4) | Agnostic | 12 (12.4) |
| Not employed, looking for work | 9 (9.3) | Information | 9 (9.3) | Atheist | 12 (12.4) | ||
| Other: please specify | 7 (7.2) | Finance and Insurance | 8 (8.2) | Catholic | 10 (10.3) | ||
| Retired | 6 (6.2) | Prefer not to answer | 6 (6.2) | Nothing in particular | 8 (8.2) | ||
| Disabled, not able to work | 5 (5.2) | Retail Trade | 6 (6.2) | Muslim | 4 (4.1) | ||
| Not employed, NOT looking for work | 2 (2.1) | Manufacturing | 5 (5.1) | Something else, Specify | 4 (4.1) | ||
| Prefer not to disclose | 1 (1.0) | Educational Services | 5 (5.1) | Buddhist | 2 (2.1) | ||
| Arts, Entertainment, and Recreation | 5 (5.1) | Hindu | 1 (1.0) | ||||
| Accommodation and Food Services | 4 (4.1) | Jewish | 1 (1.0) |
E.3. Results
To explore the possible impact of explicitly weighing harms and benefits of a use case on participant’s decision, we analyzed the participant’s judgment of acceptability before and after explicit weighing of harms and benefits (Study 2; see § LABEL:sssec:survey-qs for details on questions asked). The Type III ANOVA with Satterthwaite’s method for measurement time (before, after) indicated a marginally significant effect on usage judgment weighed by confidence, which suggests that explicit harms and benefits weighing may have an influence, albeit not at conventional significance levels. We further analyzed reasoning effect on each subset of data pertinent to each use case through a mixed effects regression model with judgment metric as a dependent variable and measurement time as an independent variable with random effect from subject. Interestingly, the result was significant for Customized Lifestyle Coach AI across different judgments including, existence (), confidence-weighed existence (), and confidence-weighed usage judgments (). Explicit weighing also had a significant effect on confidence of existence judgment for Digital Medical Advice AI (). The negative coefficients for Customized Lifestyle AI suggests that weighing harms and benefits caused participants to lower acceptance and positive coefficient to confidence on judgments on Digital Medical AI suggests that weighing harms and benefits solidified decisions. These diverging effects signify an interesting interaction between use cases and explicit weighing of harms and benefits.

### v2

@@
-Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases
-Abstract.
+Why (not) use AI? Analyzing People’s Reasoning and Conditions for AI Acceptability
+Abstract
 In recent years, there has been a growing recognition of the need to incorporate lay-people’s input into the governance and acceptability assessment of AI usage.
-However, how and why people judge different AI use cases to be acceptable or unacceptable remains under-explored.
-In this work, we investigate the attitudes and reasons that influence people’s judgments about AI’s development via a survey administered to demographically diverse participants (N=197).
-We focus on ten distinct professional (e.g., Lawyer AI) and personal (e.g., Digital Medical Advice AI) AI use cases to understand how characteristics of the use cases and the participants’ demographics affect acceptability.
+However, how and why people judge acceptability of different AI use cases remains under-explored, despite it being crucial towards understanding and addressing potential sources of disagreement.
+In this work, we investigate the demographic and reasoning factors that influence people’s judgments about AI’s development via a survey administered to demographically diverse participants (N=197).
+As a way to probe into these decision factors as well as inherent variations of perceptions across use cases, we consider ten distinct labor-replacement (e.g., Lawyer AI) and personal health (e.g., Digital Medical Advice AI) AI use cases.
 We explore the relationships between participants’ judgments and their rationales such as reasoning approaches (cost-benefit reasoning vs.
 rule-based).
-Our empirical findings reveal number of factors that influence acceptance such as general negative acceptance and higher disagreement of professional usage over personal, significant influence of demographics factors such as gender, employment, and education as well as AI literacy level, and reasoning patterns such as rule-based reasoning being used more when use case is unacceptable.
-Based on these findings, we discuss the key implications for soliciting acceptability and reasoning of AI use cases to collaboratively build consensus.
-Finally, we shed light on how future FAccT researchers and practitioners can better incorporate diverse perspectives from lay people to better develop AI that aligns with public expectations and needs.
-1. Introduction
-There are growing calls to regulate AI’s development and integration into society (Pistilli et al., 2023).
-These efforts, as reflected in the EU AI Act (Parliament, 2023), NIST AI Risk Management framework (of Standards and Technology, 2023), and recent U.S.
-Executive Order (exe, 2023), have resulted in discussions about whether certain AI use cases should be pursued at all.
-Despite much progress in this area, it is still not clear how to determine which use cases should be pursued or more heavily regulated.
-Further, little is known about how lay community members, especially those from marginalized groups, feel about the development of specific AI use cases (ada, 2023; Suresh et al., 2024).
-One significant challenge when evaluating the acceptability and impact of specific AI use cases111By use cases, we mean specific scenarios, applications, or problems that an AI system is designed to solve or assist within real-world contexts.
-is that there can be both positive and negative effects, depending on the context(Mun et al., 2024).
-For instance, while educational AI can provide affordable and accessible personal tutor, it can also lead to over-reliance of students and diminish the goal of education (Times, 2024; Zhai et al., 2024).
-To develop a generalizable approach to making decisions about AI use cases, ideally we would understand how people resolve these conflicts.
-More specifically, first, it is essential to understand differences in judgments about acceptability and likely usage across use cases, and how such judgments relate to scenario characteristics.
-Second, we need better understanding of the personal factors influencing these judgments, especially as they relate to demographic differences (Kingsley et al., 2024).
-Third, we need to better understand the reasoning strategies participants use when making judgments about AI use cases, and how those strategies do or do not relate to the judgments that are ultimately made.
-To form governance and policy decisions that anticipate and address disagreements about the development or regulation about specific AI use cases, these understandings are crucial, especially across groups of people with diverse backgrounds, experiences, and familiarity with AI.
-To address these needs, in this work we examine how and why lay people judge various AI use cases as acceptable or unacceptable, asking the following research questions:
-RQ1
-How do judgments of acceptability vary across a set of distinct AI use cases and their characteristics?
-RQ2
-What attributes or characteristics of people explain the variation in acceptability judgments?
-RQ3
-How do people reason through acceptability judgments of AI use cases?
+Our empirical findings reveal a number of factors that influence acceptance.
+We find lower acceptance of labor-replacement usage over personal health, significant influence of demographics factors such as gender, employment, education, and AI literacy level, and prevalence of rule-based reasoning for unacceptable use cases.
+Moreover, we observe unified reasoning type (e.g., cost-benefit reasoning) leading to higher agreement.
+Based on these findings, we discuss the key implications towards understanding and mitigating disagreements on the acceptability of AI use cases to collaboratively build consensus.111We will release data upon acceptance.
+1 Introduction
+There is a growing call from the public and experts alike to regulate the development and integration of AI into society (Pistilli et al.
+2023; McClain 2025).
+These efforts, as reflected in the EU AI Act (Parliament 2023), NIST AI Risk Management framework (of Standards and Technology 2023), and recent U.S.
+Executive Order (exe 2023), have resulted in discussions about whether certain AI use cases should be pursued at all.
+As these high-stakes decisions shape the future of AI, it is essential to equitably determine which AI use cases warrant pursuit despite potential harms, requiring diverse public and expert perspectives.
+Furthermore, to mitigate potential disagreements, we must understand how individuals make such decisions.
+One significant challenge when evaluating the acceptability and impact of AI use cases is that their effects can be simultaneously positive and negative, depending on the context of use, functionality, and broader societal implications (Mun et al.
+2024).
+For instance, while educational AI can provide affordable and accessible personal tutor, it can, at the same time, lead to over-reliance of students and diminish the goal of education (Times 2024; Zhai, Wibowo, and Li 2024).
+Thus, as AI is applied across increasingly diverse domains, understanding how people make decisions about its use—especially when benefits and harms conflict—becomes critical for anticipating and addressing disagreements about specific use cases.
+To tackle this challenge, we investigate the factors and reasoning that shape judgments of AI acceptability.
+First, we assess how judgments about the acceptability of development and use vary across different AI use cases,222By use cases, we mean specific real-world scenarios or problems that an AI system is designed to address.
+and how they relate to scenario characteristics (RQ1).
+Second, we explore personal factors influencing these judgments, especially as they relate to demographic differences (Kingsley et al.
+2024) (RQ2).
+Third, we analyze reasoning strategies participants use when making judgments about AI use cases, and how those strategies do or do not relate to the judgments that are ultimately made (RQ3).
 To answer these questions, we develop a survey to collect judgments and reasoning processes of 197 demographically diverse participants with varying levels of experience with AI.
 We ask participants to report whether a certain AI use case should be developed or not, whether they would use such a system, and ask them to provide rationales for their judgment and conditions that would cause them to change their judgments (Figure 1).
-We examine ten different AI use cases222We focus on text-based, non-embodied, digital systems, and while we do not specifically discuss the AI user and subject, in our use case description, we follow three of the five concepts used in EU AI Act to describe high risk use cases (Golpayegani et al., 2023): the domain, purpose, and capabilities..
-To account for differences between sectors or domains, we select use cases in two categories, professional and personal use, and vary them by required entry-level education and EU AI risk level (Table 1).
-To meaningfully differentiate and analyze participants’ reasons and reasoning strategies, we borrow concepts from moral psychology and philosophy.
-We investigate participants’ rationales through two distinct but sometimes overlapping reasoning patterns: cost-benefit reasoning, which assesses expected outcomes (e.g., ”using AI for this task would save time”), and rule-based reasoning, which evaluates the intrinsic values of the action itself (e.g., ”having humans perform this task would be inherently wrong”) (Cushman, 2013; Cheung et al., 2024).
-We further explore the moral foundations reflected in participants’ reasoning, with moral foundations theory333We used the five foundational dimensions: Care, Fairness, Loyalty, Authority, and Purity.
-Although these dimensions have been updated to encompass a broader range of values beyond WEIRD (White, Educated, Industrialized, Rich, and Democratic) populations (Atari et al., 2023), we selected this version for survey brevity.
-(Graham et al., 2011, 2008).
-Additionally, to understand aspects of AI that raise concerns, we employ three dimensions based on prior studies (Solaiman et al., 2023; Mun et al., 2024): functionality (system capabilities like performance, bias, and privacy), usage (context of system integration, such as supervision, misuse, or unintended use), and societal impact (effects on individuals, communities, and society, such as job loss and over-reliance).
-Our empirical results show general higher acceptance of personal use cases over professional.
-While both categories of use cases show decreased acceptance with increased entry level education and risk, professional use cases display more variability and disagreements across judgments (RQ1).
+We perform a focused investigation of acceptability using ten AI use cases333We focus on text-based, non-embodied, digital systems, and while we do not specifically discuss the AI user and subject, in our use case description, we follow three of the five concepts used in EU AI Act to describe high risk use cases (Golpayegani, Pandit, and Lewis 2023): the domain, purpose, and capabilities.
+that we systematically select for different risk levels, spanning two highly-discussed domains with ongoing efforts to develop such use cases: personal health and labor replacement (McClain 2025; Kelly 2025; Kolata 2024; Pierson et al.
+2025; Rajpurkar et al.
+2022; Lee 2024).
+To understand characteristics of AI use cases that might affect perceptions beyond category, we vary them by required entry-level education and EU AI risk level (Table 1).
+We perform a multi-pronged analyses of people’s rationales.
+Drawing from moral philosophy, we examine participants’ answers using two reasoning patterns: cost-benefit reasoning, which assesses expected outcomes (e.g., “using AI for this task would save time”; akin to utilitarian reasoning), and rule-based reasoning, which evaluates the intrinsic values of the action itself (e.g., “having humans/AI perform this task would be inherently wrong”; akin to deontological reasoning) (Cushman 2013; Cheung, Maier, and Lieder 2024).
+We then analyze the moral frameworks participants apply, drawing on moral foundations theory (Graham et al.
+2011, 2008), to identify the dimensions they prioritize in decision making.
+Finally, to understand conditions under which participants might flip there decisions, we employ three dimensions based on prior studies (Solaiman et al.
+2023; Mun et al.
+2024): functionality (system capabilities like performance, bias, and privacy), usage (context of system integration, such as supervision, misuse, or unintended use), and societal impact (effects on individuals, communities, and society, such as job loss and over-reliance).
+Our empirical results show general higher acceptance of personal health use cases over labor-replacement.
+While participants’ acceptability judgments decreased with increased entry-level education and risk for each category respectively, professional use cases display more variability and disagreements across judgments (RQ1).
 Acceptability significantly varied among demographic groups and levels of AI literacy, with lower acceptability observed particularly among non-male participants and those familiar with AI ethics (RQ2).
-Finally, our results show varying distribution of reasoning types across acceptability decisions with rule-based reasoning being associated with negative acceptance as well as concern for societal impact.
-Further qualitative analysis reveal rules such as the need for humanness in certain use cases whether it be for empathy or interaction (RQ3).
+Finally, our results show varying distribution of reasoning types across acceptability decisions, with rule-based reasoning being associated with negative acceptance and unified reasoning types showing higher agreement.
+Further qualitative analysis revealed participants’ normative assumptions about AI, humanness, and society—for example, viewing empathy as essential to humanness but lacking in AI (RQ3).
 Our findings shed novel light onto the diversity of people’s acceptability and reasoning of AI uses in distinct domains and risk levels.
-We conclude with a discussion highlighting three key implications: first, diverse methodologies are needed to effectively analyze use cases and their characteristics; second, involving diverse stakeholders is crucial for assessing the acceptability of AI applications, particularly in workplaces; and third, further investigation into human reasoning processes about AI, notably rule-based reasoning, is needed to inform consensus-building in policy making.
-2. Related Works
-In this section, we briefly summarize the background and related work towards assessing acceptability and impact of AI use cases. In each subsection, we highlight how our work extends prior work.
-2.1. Assessing Impact of AI
-Recent years have witnessed increasing calls from academics (Kieslich et al., 2023; Hecht et al., 2021; Bernstein et al., 2021; Ashurst et al., 2020; Lin et al., 2020; CVPR, Ethics Guidelines, 2023; ACL, Ethic Policy, 2023; ICML, Publication Ethics, 2023; Olteanu et al., 2023; Center for Advanced Study in the Behavioral Sciences, 2020), government (The Ada Lovelace Insitute, 2022; The White House, 2023; National Artificial Intelligence Research Resource Task Force, 2023), civil society (on AI, 2021; Ada Lovelace Institute, 2022; The Ada Lovelace Insitute, 2022; Society, 2023; Metcalf et al., 2021; Reisman et al., 2018), and industry (Microsoft, 2022a, b; Walker and Croak, 2021; OpenAI, 2022; Hecht et al., 2021; Deng et al., 2024) to assess the impact of AI systems designed and developed by AI researchers and practitioners.
-This effort has particularly highlighted the need to understand the positive impact while grappling with the potential negative impact of integrating AI into certain products and services that affect people’s daily lives (Ashurst et al., 2020; Lin et al., 2020; Hecht et al., 2021).
-In response, researchers in FAccT, HCI, and AI have developed tools and processes to support AI researchers and practitioners in anticipating the impact of AI systems they developed(Wang et al., 2024; Kieslich et al., 2023; Buçinca et al., 2023; Deng et al., 2024; Weidinger et al., 2022).
-For example, many have developed AI impact taxonomies or checklists to help developers categorize AI impact(Weidinger et al., 2022; Microsoft, 2022b).
-Wang et al.
-and Deng et al.
-developed tools and templates to support industry AI developers and researchers in assessing the potential negative societal impact of their work, such as job displacement or stereotyping social groups.
-However, this prior work primarily focuses on supporting AI experts rather than diverse lay people’s impact assessments of potential AI use cases.
-Our work extends these prior efforts by understanding diverse (and sometimes conflicting) perspectives on both positive and negative impact of AI use cases from lay people, as a crucial step to complement the AI impact assessments conducted by AI researchers and practitioners.
-2.2. Understanding People’s Perceptions of AI Use Cases
-Responding to the calls on meaningfully engaging lay people in assessing the impact of specific AI use cases, prior work have started to understand lay people’s perceptions of AI use case (Buçinca et al., 2023; Kieslich et al., 2023; Mun et al., 2024; Kingsley et al., 2024).
-Among other findings, this prior work revealed a substantial amount of disagreement regarding the desired behavior of AI, primarily due to the subjectivity inherent in certain tasks (e.g., toxicity detection (Sap et al., 2019; Blodgett et al., 2020), image captioning (Zhao et al., 2021)) and ambiguous ethical implication of decisions made by AI for certain tasks (e.g., self driving cars (Awad et al., 2018), medical AI (Chen et al., 2023), predictive analysis (Barocas and Selbst, 2016)).
-Work done by Mun et al.
-highlighted that lay people can envision diverse set of harms specific to different AI use cases, complementary to those defined by experts.
-Another line of work also begins to examine how factors such as demographic backgrounds and previous exposure to discrimination can affect people’s sensitivity towards potential AI harms (Kingsley et al., 2024).
-Our work extends this prior work by examined the detailed reasoning processes of lay people regarding the acceptability and the trade-offs between positive and negative impacts of AI use cases.
-In particular, we draw on model decision theory framework, such as the moral foundations developed by Graham et al., to design a survey flow (See Figure 1 to solicit lay people’s decision-making processes (and potential moral conflicts) when assessing both the benefits and harms of concrete AI use cases.
-2.3. Background: Moral Decision Making
-Morality, characterized by diverse values across cultures and social groups, aims to suppress selfishness to facilitate social life (Kesebir and Haidt, 2010).
-To understand decision-making in AI use cases, we draw on moral psychology and dual system theory.
-We examine two decision-making systems: cost-benefit reasoning, which assesses outcomes and consequences, and rule-based reasoning, focusing on norms, rules, and virtues (Cushman, 2013; Cheung et al., 2024).
-These correspond to utilitarian reasoning (maximizing good) and deontological reasoning (duties and rights), respectively.
-Additionally, we apply moral foundations theory (Graham et al., 2008) to identify values and potential moral conflicts in AI development.
-3. Study Design and Data Collection
-To understand the decision making factors and reasoning patterns of a diverse population regarding AI use cases, we conducted a survey-based study with demographically diverse participants.
+We conclude with a discussion highlighting three key implications for future researchers, practitioners, and policymakers working on advancing ethical and responsible AI development: first, diverse methodologies are needed to effectively analyze use cases and their characteristics; second, involving diverse stakeholders is crucial for assessing the acceptability of AI applications, particularly in workplaces; and third, further investigation into human reasoning processes about AI, notably rule-based reasoning, is needed to inform consensus-building in policy making.
+2 Related Works
+While there were many efforts towards ethical AI development and deployment by academics (Kieslich, Diakopoulos, and Helberger 2023; Bernstein et al.
+2021; Lin et al.
+2020), industry (OpenAI 2022; Deng, Barocas, and Vaughan 2024), and government (The White House 2023), they have largely lacked diverse public inputs.
+To address this gap, many works from both academia and civil society have sought to meaningfully engaging lay people in assessing the impact of specific AI use cases.
+Prior works have focused on anticipating harms (Buçinca et al.
+2023) and impacts (Kieslich, Diakopoulos, and Helberger 2023; ada 2023) through participatory foresight, uncovering diverse, sometimes diverging, viewpoints about AI biases and values (Kingsley et al.
+2024; Jakesch et al.
+2022; Kapania et al.
+2022), and governance efforts of AI (Zhang and Dafoe 2020).
+However, to the best of our knowledge, only Mun et al.
+considered development decisions by diverse lay-users with an option of not developing a use cases.
+Among other findings, these prior works from AIES and broader Responsible AI venues revealed a substantial amount of variations in perceptions primarily among demographic lines (e.g., gender, race, political leaning) regarding the desired behavior of AI.
+However, little attention has been given to identifying reasoning of participants over AI use cases.
+While some works have identified decision variations under ambiguous ethical implications of decisions made by AI for certain tasks (e.g., self driving cars (Awad et al.
+2018), medical AI (Chen et al.
+2023), predictive analysis (Barocas and Selbst 2016)) and inherent value conflicts (Jakesch et al.
+2022), these works have not focused on self-reported reasoning.
+Thus, our work addresses gap by closely examining the detailed, self-reported reasoning processes of lay people regarding the acceptability of AI use cases without explicitly guiding towards outcome-based (i.e., utilitarian) or value-based (i.e., deontological) reasoning, allowing participants to freely choose and express their deliberation process.
+3 Study Design and Data Collection
+To answer our research questions on how and why people judge AI use cases as acceptable, we conducted a survey-based study with demographically diverse participants.
 In this section, we discuss the selection of use cases (§ 3.1), survey design (§ 3.2), and data collection details and participant demographics (§ 3.3).
-3.1. Use Cases
-To understand how different characteristics of AI use cases can impact judgments and decision making processes, we crafted ten different AI use cases.
-We first chose two broad application categories frequently mentioned by the public in previous works (Kieslich et al., 2024; Mun et al., 2024): AI in personal, everyday usage where participants could uniformly consider themselves as AI users and AI in professional usage where AI takes on a role thus far done by a human as a profession, e.g., Lawyer AI.
-We then developed five use cases in the AI in professional usage category that varied according to the level of education required for entry and five use cases in AI in personal category that varied according to the risk level assigned by the EU AI act.
-All the personal use cases were in the health domain, reflecting a key domain of interest conveyed by lay users in previous works (Mun et al., 2024; Kieslich et al., 2024).
+3.1 Use Cases
+To answer RQ1 which examines the impact of different characteristics AI use cases on judgments and decision making processes, we carefully crafted ten different AI use cases as vignettes.
+We first chose two broad application categories frequently mentioned by the public in previous works (Kieslich, Helberger, and Diakopoulos 2024; Mun et al.
+2024): AI in labor-replacement where AI takes on a role in society thus far done by a human as a profession (e.g., Lawyer AI), and AI in personal health, where participants could uniformly consider themselves as AI users.
+We systematically developed five use cases for each category, varying by required education level for labor-replacement applications and by EU AI Act-assigned risk level for personal health applications.
 | Use Case | Factor | Description |
-| Professional Use Cases | ||
-| Lawyer | Doctoral/Professional Degree | Advises clients on digital legal proceedings/transactions. |
-| Elementary School Teacher | Bachelor’s degree | Teaches academic skills at the elementary school level. |
-| IT Support Specialist | Some college, no degree | Maintains computer networks and provides technical help. |
-| Government Eligibility Interviewer | High school diploma | Determine eligibility for government programs/resources. |
-| Telemarketer | No formal education | Solicits donations or orders over the telephone. |
-| Personal Use Cases | ||
-| Digital Medical Advice | High Risk | Provide medical assessments prior to medical consultations. |
-| Customized Lifestyle Coach | High / Limited Risk | Personalized advice for healthy living and wellness. |
-| Personal Health Research | Limited Risk | Summarizes research related to personal health issues. |
-| Nutrition Optimizer | Limited / Low Risk | Personalize meals and optimize nutritional intake. |
-| Flavorful Swaps | Low Risk | Suggest delicious and healthy alternatives food options. |
-Professional Use Case Scenarios
+| Labor-replacement Use | ||
+| Lawyer | Doctoral/Prof | Digital legal advice |
+| Elementary Teacher | Bachelor | Teaches elementary students |
+| IT Support | Some college | Maintains networks; tech support |
+| Eligibility Interviewer | High school | Determines benefit eligibility |
+| Telemarketer | None | Calls to sell/solicit |
+| Peronal Health Use | ||
+| Digital Medical Advice | High Risk | Medical advice pre-consultation |
+| Lifestyle Coach | High/Limited | Personalized wellness advice |
+| Health Research | Limited | Summarizes personal health info |
+| Nutrition Optimizer | Lim./Low | Personalized meal/nutrition tips |
+| Flavorful Swaps | Low | Suggests healthy food alternatives |
+Labor-replacement Use Case Scenarios
 For the first area of focus, AI in labor replacement, we collected jobs listed in the U.S.
 census bureau444https://www.bls.gov/ooh/occupation-finder.htm and sorted them according to entry level education required as stated in the census.
-We chose education level as it has been tightly linked to socioeconomic and occupational status (Svensson, 2006; Evetts, 2006).
+We chose education level as it has been tightly linked to socioeconomic and occupational status (Svensson 2006; Evetts 2006).
 We selected jobs that have a large portion of digital or intellectual components with minimal requirement for embodiment resulting in following five professional roles: Lawyer, Elementary school teacher, IT support specialist, Government support eligibility interviewer, and Telemarketer.
 See Table 1 for further details.
-Personal Use Case Scenarios
-To understand the acceptability of different health applications in personal and private life, we adapted the descriptions of personal use cases written by participants from prior works (Mun et al., 2024; Kieslich et al., 2024).
-The research team iteratively refined the descriptions to reflect risk levels according to EU AI Act, and confirmed agreement with categories assigned by GPT-4, following Herdel et al..
+Personal Health Use Case Scenarios
+To understand the acceptability of different health applications in personal and private life, we drew from use cases written by participants from prior works (Mun et al.
+2024; Kieslich, Helberger, and Diakopoulos 2024) to systematically craft use cases which varied by risk levels according to EU AI Act.
+We ensured accurate reflection of the risk levels through iterative refinement of descriptions and agreement with categories assigned by GPT-4, following Herdel et al..
 See Table 1 for further details.
-3.2. Survey Design
-Our survey presents participants with five use case descriptions, all from either the professional or personal category (randomly assigned and presented in random order) (see § 3.1 for details).
-After each description, participants answer: ”Do you think a technology like this should be developed?” (Q1) and then, ”How confident are you in your above answer?” (Q2).
-They provide open-text rationales by completing the prompt, ”[Use Case] should be developed because…” (Q3), tailored to their Q1 response (”should” for ”Yes” and ”should not” for ”No”).
-Participants also specify a condition for switching their decision with, ”[Use Case] should not be developed if…” (Q4), adjusted similarly to Q1.
-Subsequently, they answer, ”If [Use Case] existed, would you use its service?” (Q5) and express confidence with, ”How confident are you in your above answer?” (Q6).
+3.2 Survey Design
+Our survey presents participants with five use case descriptions in random order, all from randomly assigned category, labor-replacement or personal health (see § 3.1 for details).
+After each description, participants answer: “Do you think a technology like this should be developed?” (Q1) and then, “How confident are you in your above answer?” (Q2).
+To allow for examining of their reasoning, participants then provide open-text rationales by finishing the sentence, “[Use Case] should [not] be developed because…” (Q3), adjusted dynamically depending on their answer to Q1.
+Participants also described that they would switch their opinion on acceptability of development of the use case: however, “[Use Case] should [not] be developed if…” (Q4; also dynamically rephrased based on Q1 answer).
+Subsequently, they answer, “If [Use Case] existed, would you use its service?” (Q5) and express confidence with, “How confident are you in your above answer?” (Q6).
 Refer to Table 6 in the Appendix for the exact wording of the questions.
 Collecting Participant Characteristics
-Following the main survey, we asked participants questions about their AI literacy and demographics to explore various factors affecting perception of AI acceptance.
-We adopted a shortened version of AI literacy questionnaires from previous works (Wang et al., 2023; Mun et al., 2024) with four AI literacy aspects, AI awareness, usage, evaluation, and ethics, and two additional questions for generative AI, usage frequency and familiarity with limitations.
-We collected demographic information of the participants such as race, gender, age, sexual orientation, religion, employment status, income, and level of education.
-Additionally, we collected information about chronicity, i.e., prolonged experiences of everyday discrimination, of their discrimination experiences (if any) following Kingsley et al..
-3.3. Data Collection and Participant Demographics
+Following the main survey, we asked participants questions about their AI literacy level and demographics to explore various factors affecting perception of AI acceptance (RQ2).
+We adopted a shortened version of AI literacy questionnaires from previous works (Wang, Rau, and Yuan 2023; Mun et al.
+2024) with four AI literacy aspects, AI awareness, usage, evaluation, and ethics, and two additional questions for generative AI, usage frequency and familiarity with limitations.
+We collected demographic information of the participants such as race, gender, age, sexual orientation, religion, employment status, income, and level of education; see Appendix 7 for detailed list of questions.
+Additionally, we collected information about discrimination chronicity, i.e., prolonged experiences of everyday discrimination, of their discrimination experiences (if any) following Kingsley et al..
+3.3 Data Collection and Participant Demographics
 We used Prolific555https://www.prolific.com to recruit participants.
-To represent diverse sample, we stratified our recruitment by the ethnicity categories (White, Mixed, Asian, Black and Other) and age (18-48, 49-100) provided by Prolific.
+To represent diverse sample, we stratified our recruitment by the ethnicity categories (White, Mixed, Asian, Black, and Other) and age (18-48, 49-100) as provided by Prolific.
 We also added criteria for quality such as survey approval rating and number of previous surveys completed.
@@
 See Appendix A.3 for further details on participants.
-4. Analysis Methods
+4 Acceptability & Reasoning
+Analysis Methods
 Our surveys consisted of both multiple choice (numerical) and open-text questions designed to answer our research questions.
 In this section, we detail our process for numerical (§ 4.1) and open-text (§ 4.2) analysis.
-4.1. Multiple Choice Analysis
+4.1 Multiple Choice Analysis
 We analyzed the judgment and confidence ratings by mapping judgment (Q1, Q5) to 1 (“Should be developed”, “Would use”) or -1 (“Should not be developed”, “Would not use”) and confidence (Q2, Q6) to a scale from 1 to 5.
@@
 We included a subject-specific random effect when using ANOVA and regression models and added a use-case-specific random effect when applicable.
-We factorized demographic responses for analysis with the exception of discrimination chronicity, which we aggregated to a numerical value (Kingsley et al., 2024; Michaels et al., 2019).
+We factorized demographic responses for analysis with the exception of discrimination chronicity, which we aggregated to a numerical value (Kingsley et al.
+2024; Michaels et al.
+2019).
 We also converted responses to AI literacy questions to numerical values for analysis.
-4.2. Open-response Analysis
-To assess the reasoning methods used by the participants, we analyzed the open-text responses on elaborations to their decisions (Q3) and circumstances in which their decisions would switch (Q4) along the following three dimensions: reasoning types (cost-benefit, rule-based, both, unclear), reference to moral foundations (Care, Fairness, Purity, Authority, Loyalty), and switching conditions (Functionality, Usage, Societal Impact).
-By analyzing reasoning types and moral values reflected in the participants’ justifications, we aim to characterize how participants made their decisions, and by analyzing various factors such as primary concerns and stakeholders, we aim to discover what aspects were salient for the participants in their decisions.
+4.2 Open-response Analysis
+Background: Moral Decision Making
+To understand decision-making in AI use cases, we draw on moral psychology and dual system theory.
+We examine two decision-making systems: cost-benefit reasoning, which assesses outcomes and consequences, and rule-based reasoning, focusing on norms, rules, and virtues (Cushman 2013; Cheung, Maier, and Lieder 2024).
+These correspond to utilitarian reasoning (maximizing good) and deontological reasoning (adhering to moral duties and rights), respectively.
+Additionally, we apply moral foundations theory (Graham et al.
+2008) to identify values and potential moral conflicts in AI development.
+To assess the reasoning methods used by the participants, we analyzed the open-text responses on elaborations to their decisions (Q3) and circumstances in which their decisions would switch (Q4) along the following three dimensions: reasoning types (cost-benefit, rule-based, both, unclear), reference to moral foundations666We used the five foundational dimensions: Care, Fairness, Loyalty, Authority, and Purity.
+Although these dimensions have been updated to encompass a broader range of values beyond WEIRD (White, Educated, Industrialized, Rich, and Democratic) populations (Atari et al.
+2023), we selected this version for survey brevity.
+(Care, Fairness, Purity, Authority, Loyalty), and switching conditions (Functionality, Usage, Societal Impact).
+By analyzing reasoning types and moral values reflected in the participants’ justifications, we aim to characterize how participants made their decisions, and by analyzing various factors such as primary concerns in switching condition, we aim to discover what aspects were salient for the participants in their decisions.
 Classification and Aggregation
-We classified participants’ responses to Q3 (elaboration of judgment) and Q4 (conditions for switching decisions) using OpenAI’s o1-mini666o1-mini-2024-09-12.
-To validate the model’s classification performance, results were compared with a reference set of 100 samples annotated by three independent annotators, comprised of team members and a professional annotator.
+We classified participants’ responses to Q3 (elaboration of judgment) and Q4 (conditions for switching decisions), totaling 985 samples for each question, using OpenAI’s gpt-4o777gpt-4o-2024-11-20.
+To validate the model’s classification performance, results were compared with a reference set of 100 samples annotated by three independent annotators, comprised by members of the research team and a professional annotator.
 Initially, each annotator independently assessed the data, and then consensus was reached through discussion to establish a gold standard set.
-The inter-rater agreement between the gold standard and o1-mini’s annotations was evaluated using Gwet’s AC1 metric, chosen for its robustness with infrequent labels (Wongpakaran et al., 2013).
-The agreement levels varied, ranging from moderate to almost perfect with no dimension below moderate.
-Annotations for cost-benefit reasoning, rule-based reasoning, and authority reached near-perfect agreement; purity and usage achieved substantial agreement; the rest showed moderate agreement.
-Due to a lack of sufficient test samples and minimal occurrences in the annotated data, the moral value dimension Loyalty was excluded from further analysis.
-The annotations were converted into a one-hot encoding format for statistical analysis.
-See Appendix B for further details.
-5. Findings
+The inter-rater agreement between the gold standard and o1-mini’s annotations was evaluated using Gwet’s AC1 metric, chosen for its robustness with infrequent labels (Wongpakaran et al.
+2013).
+While the agreement levels varied, ranging from almost perfect to moderate (0.98–0.57), all dimensions had above substantial agreement except Societal Impact.
+Annotations for cost-benefit reasoning, rule-based reasoning, and authority reached near-perfect agreement.
+Due to minimal occurrences in both human and LLM annotations, the moral foundation dimension Loyalty was excluded from further analysis.
+The annotations were conducted based on presence or absence of the values and were converted into binary format for statistical analysis.
+See Appendix C for further details on agreement and automatic annotation settings.
+5 Findings
 Our work aims to uncover variations in acceptability of AI use cases and factors and reasoning processes that underlie these judgments.
 In this section, we discuss our findings about the judgments of the AI use cases (§5.1), personal factors that may influence the decision such as demographics and AI literacy (§5.2), and factors in rationales that could uncover reasoning processes that lead to judgments (§5.3).
-5.1. RQ1. Perceptions and Disagreements of Use Cases
-5.1.1. Use Case Factors
-In our analysis, we investigated the effects of use cases on participants’ judgments of two categories—professional and personal—along with ten specific use cases on judgments and confidence about their development and usage.
-Figure 2 illustrates significant differences in acceptability depending on the categories.
-Notably, personal use cases had higher acceptability (; ) than professional use cases (; ).
-These differences are more pronounced when judgments are weighted by confidence.
-Although levels of confidence did not differ significantly between categories, they were fairly high for both categories, and personal category exhibited slightly lower confidence for usage.
-Professional Use Cases
-Exploring specific use cases within the professional category, we observed that Elementary School Teacher AI (; ) had the lowest acceptability for both types of judgments followed by Lawyer AI () and Telemarketer AI ().
-Interestingly, IT Support Specialist AI had the highest acceptability (; ).
-While all other use cases showed higher acceptability for usage over development, Telemarketer AI uniquely had higher acceptance for development over usage (; ).
-Additionally, confidence on using Telemarketer AI () was significantly () higher than that of Government Eligibility Interviewer AI (), which had the lowest confidence in usage.
-Near 0 mean for development of Lawyer AI () and usage for Telemarketer AI () suggest disagreement within judgments.
-Moreover, Elementary School Teacher AI uniquely unacceptable across both acceptability judgments (, ) underscoring a unique characteristic perhaps related to care.
-See Figure 3 for further details
-Personal Use Cases
-In personal use scenarios, Digital Medical Advice AI (; ), reflecting high risk level, consistently had lower acceptance across judgment types, compared to all other use cases.
-However, Customized Lifestyle Coach AI had the lowest confidence across both judgments (; ).
-Nutrition Optimizer (; ) had the highest mean acceptance across the two acceptability judgments.
-Interestingly, unlike the professional use cases, which had slightly higher acceptance for usage, personal use cases had lower acceptance for usage in general compared to development.
-Notably, many personal use cases had substantially high acceptance, especially compared to professional use cases.
+5.1 RQ1. Use Case Perceptions & Disagreements
+In our analysis, we investigated the effects of use cases on participants’ judgments using our ten use case vignettes.
+Overall, acceptability statistically differed among the two categories (; ).
+Notably, personal health use cases had higher acceptability (; ) than labor-replacement use cases (; ).
+See Figure 7 in Appendix B for additional category comparison results.
+Labor-replacement Use Cases
+Exploring specific use cases within the labor-replacement category (Figure 2), we observed that Elementary School Teacher AI (; ) had the lowest acceptability for both types of judgments followed by Lawyer AI () and Telemarketer AI () where their near zero mean suggest disagreement within judgments.
+Interestingly, IT Support Specialist AI had the highest acceptability (; ) despite human replacement potential and median education level.
+Personal Health Use Cases
+In personal health use scenarios, Digital Medical Advice AI (; ), reflecting high risk usage, consistently had lower acceptance across judgment types, compared to all other use cases.
+Nutrition Optimizer (; ) had the highest mean acceptance across both acceptability judgments.
+Interestingly, unlike the labor-replacement use cases which had slightly higher acceptance for usage, personal use cases had lower acceptance for usage in general compared to development.
 | DEV() | USAGE() | |||
-| Use Case Variation | Judgment | Confidence | Judgment | Confidence |
-| Professional | ||||
-| Education Level | ||||
-| Intercept | ||||
-| Personal | ||||
-| EU AI Risk Level | ||||
-| Intercept | ||||
-5.1.2. Use Case Variations
-When selecting use cases, we used two underlying variations: entry level of education required for professional use cases and EU AI risk levels for personal use cases.
-As risk levels and required education increased, we observe consistent negative effects on judgments, with personal use cases (; ) showing stronger effects compared to professional scenarios (), where only development judgments were significantly associated.
-Again, acceptance for personal use cases (; ) were higher than professional use cases ( ) as conveyed by the intercepts.
-Confidence ratings remained consistently high across all conditions (intercepts ¿ 3.97), though they showed a small but significant decrease with increasing risk levels in personal use cases (; ), while professional use cases showed no significant impact on confidence.
-Thus, while education level required for entry does show some significant effect, the EU AI risk level had consistent effects with higher significance.
-5.1.3. Disagreements
-Observing the standard deviation and visualization of the distribution shows further understanding of possible disagreements among use cases.
-We compare the judgments weighted by confidence to understand not only the differences in judgment but also their strength.
-Interestingly, the use cases with four highest disagreements in both judgments were all professional uses in order of Telemarketer (; ), Elementary School Teacher (; ), Lawyer (; ), and Government Eligibility Interviewer AI (; ).
+| Judg. | Conf. | Judg. | Conf. | |
+| Labor-replacement | ||||
+| Coeff. | ||||
+| Personal Health | ||||
+| Coeff. | ||||
+Use Case Variations
+When selecting use cases, we used two underlying variations: entry level of education required for labor-replacement use cases and EU AI risk levels for personal health.
+As risk levels and required education increased, we observe consistent negative effects on judgments, with personal health use cases showing stronger effects (; ) compared to labor-replacement scenarios (), where only development judgments were significantly associated.
+Confidence ratings showed a small but significant decrease with increasing risk levels in personal health use cases (; ), while labor-replacement use cases showed no significant impact on confidence.
+Disagreements
+We compare the standard deviation of judgments weighted by confidence to understand possible disagreements and their strength among use cases.
+Interestingly, the use cases with four highest disagreements in both judgments were all labor-replacement uses in order of Telemarketer (; ), Elementary School Teacher (; ), Lawyer (; ), and Government Eligibility Interviewer AI (; ).
 These four use cases were followed by Digital Medical Advice AI (, ).
 The use cases with the lowest disagreements were surprisingly Nutrition Optimizer (; ) followed by IT Support Specialist AI (; ).
-These results underscore the general controversy of professional usages but also shows that acceptability is highly use case dependent.
-5.2. RQ2. Impact of Personal Factors on Acceptability Judgment
-5.2.1. Demographic Factors
+RQ1 Takeaways
+Our results show that there are significant variation in judgments based on use case characteristics such as category, EU-defined risk level, and, to some extent, required education level for labor-replacement cases.
+The uniquely negative response to the Elementary School Teacher AI highlights potential concerns specific to care work.
+High disagreement in labor-replacement scenarios underscores the need for cautious integration of AI into existing roles.
+In contrast, the consistent positive judgments for IT Support Specialist AI suggest that not all labor-replacement use cases are viewed equally, indicating the need for nuanced understandings of acceptability.
+We further examine this variability in § 5.3.
 | DEV ( (SE)) | USAGE ( (SE)) | |||||
@@
 | High | ||||||
-In analyzing the demographic factors influencing judgments on the development and usage of AI use cases, several significant trends emerged from the data.
-First, we observed that the demographic variables had less than 0.5 correlation, except for age 65+ and Retired employment status.
-Across both categories of use cases, age was positively associated with confidence in acceptability judgment of usage for age 25-34 (, ) and 55-64 (, ).
-Race also had notable influences; specifically, Asian participants exhibited significantly lower confidence in both development and usage judgments (, ; , ), particularly in professional contexts.
-Gender emerged as a crucial determinant, with non-male participants consistently showing negative judgments across both development (, ) and usage (, ), indicating potential discrepancies in perception or experience with AI applications.
+5.2 RQ2. Impact of Personal Factors on Acceptability Judgment
+Demographic Factors
+As shown in Table 3, several demographic factors significantly influenced use case judgments.888Prior to analysis, we observed that the independent variables in the model had less than 0.5 correlation, except for age 65+ and Retired employment status.
+Across both categories of use cases, certain age groups were positively associated with confidence in usage: 25-34 (, ) and 55-64 (, ).
+Race also had notable influences; specifically, Asian participants exhibited significantly lower confidence in both development and usage judgments (, ; , ), particularly in labor-replacement contexts.
+Gender emerged as a crucial determinant, with non-male participants consistently showing negative judgments (, ; , ), indicating potential discrepancies in perception or experience with AI applications.
 Liberal views, especially among those identifying as strongly liberal, were associated with negative judgments across both categories of use cases (, ; , ), suggesting a skeptical stance towards AI’s prevalence and role.
-Employment hours also contributed, with individuals working 40+ hours per week displaying a positive association with development judgments (, ), suggesting more exposure or reliance on AI use cases.
-High experience of discrimination chronicity was significantly related to lower acceptance of development, , .
-These findings highlight the significant role of demographic factors in shaping perceptions and attitudes toward AI technologies.
+Employment hours also contributed, with individuals working 40+ hours per week displaying a positive association with development (, ), suggesting more exposure or reliance on AI use cases.
+High experience of discrimination chronicity was significantly related to lower acceptance of development (, ).
 See Table 25 in the Appendix for ANOVA results.
@@
 | AI Literacy | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
-| Professional | ||||||
+| Labor-replacement | ||||||
 | AI Ethics | ||||||
@@
 | Gen AI Limit. Familiarity | ||||||
-| Personal | ||||||
+| Personal Health | ||||||
 | AI Skills | ||||||
@@
 | Gen AI Limit. Familiarity | ||||||
-5.2.2. AI Literacy
+AI Literacy
 We identified a correlation greater than 0.5 among three AI literacy aspects: awareness, usage, and evaluation.
-These were aggregated into a single factor, AI Skills.
-As shown in Table 4, understanding of AI Ethics was associated with lower acceptability for both personal (, ; , ) and professional ().
-However, across both categories of use cases, high Generative AI Usage Frequency resulted in higher acceptance for both professional (, ; , ) and usage acceptance (, ; , ).
-Notably, for personal use cases, AI Skills was positively associated with confidence of judgments (, for both development and usage) and judgment weighted by confidence (, ), while Generative AI Limitation Familiarity was positively associated with confidence for professional usage (, ; , for usage).
-These results implicate that different understandings of and experiences with AI can impact judgments of acceptability, corroborating previous findings (Kramer et al., 2018).
-5.3. RQ3. Factors in Participant Rationale
-To provide deeper insights to our analyses of acceptability judgments, we also examined open-text rationales (Q3) for judgments of development (Q1) and conditions for switching their decisions (Q4).
-5.3.1. Decision-making Types
-As we defined in § 4.2, we focus on two distinct reasoning types for decision-making, distinguished by consideration of outcome versus consideration of value inherent in action: cost-benefit reasoning (e.g., “it gives more people access to medical advice and treatment”, P365) and rule-based reasoning (e.g., should not be developed because “human interaction is better”, P249).
-As shown in Figure 5, generally, participants used more cost-benefit reasoning with highest percentage for IT Support Specialist (91.0%) and Nutrition Optimizer (92.8%).
-This result is particularly interesting as we observed these two use cases to have the lowest disagreement (see § 5.1.1) and signifies that perhaps unified reasoning type leads to less disagreement.
-On the other hand, rationales for Elementary School Teacher AI contained most percentage of rule-based reasoning (30.0%) followed by Lawyer AI (22.0%).
-Interestingly, Elementary School Teacher AI and Lawyer AI were the use cases with the lowest development acceptability.
-These results suggest that less acceptable AI use cases might trigger more rule-based reasoning, which could, however, be due to the lack of consensus and rules around AI, especially those that are positive.
-this suggests the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable.
-These results suggest that the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable.
-5.3.2. Moral Foundations
+To avoid multicollinearity, we aggregated them into a single factor, AI Skills.
+As shown in Table 4, understanding of AI Ethics was associated with lower acceptability for both personal health (, ; , ) and labor-replacement ().
+However, across both categories, high Generative AI Usage Frequency resulted in higher acceptance (Labor-replacement – , ; , , Personal Health – , ; , ).
+Notably, for personal health use cases, AI Skills was positively associated with confidence of judgments (, ), while Generative AI Limitation Familiarity was positively associated with confidence for labor-replacement usage (, ; , ).
+RQ2 Takeaways
+Our findings highlight the significant role of lived experiences and backgrounds, such as age, race, gender, political view, employment, and discrimination experience, in shaping perceptions and attitudes toward AI technologies.
+Moreover, our results indicate that different understandings of and experiences with AI can impact judgments of acceptability, corroborating previous findings (Kramer et al.
+2018).
+| DEV ( (SE)) | USAGE ( (SE)) | ||||||
+| Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. | ||
+| Reasoning Type | |||||||
+| Cost-benefit | |||||||
+| Rule-based | |||||||
+| Moral Value | |||||||
+| Care | |||||||
+| Fairness | |||||||
+| Authority | |||||||
+| Switching Condition | |||||||
+| Usage | |||||||
+| Societal Impact | |||||||
+5.3 RQ3. Factors in Participant Rationale
+To deepen our analysis of use case acceptability, we examined open-text rationales (Q3) and decision-switching conditions (Q4) related to development judgments (Q1).
+Decision-making Types
+As we defined in § 4.2, we focus on two distinct reasoning types for decision-making: cost-benefit reasoning, which emphasizes outcomes (e.g., “it gives more people access to medical advice and treatment”, P365), and rule-based reasoning, which reflects values inherent in the action itself (e.g., should not be developed because “human interaction is better”, P249).
+As shown in Figure 4, generally, participants used more cost-benefit reasoning, especially for the IT Support Specialist (91.0%) and Nutrition Optimizer (92.8%) use cases.
+This result is particularly interesting as we observed these two use cases to have the lowest disagreement (see § 5.1), suggesting unified reasoning type may lead to more consistent judgments.
+On the other hand, rationales for Elementary School Teacher AI contained most percentage of rule-based reasoning (30.0%) followed by Lawyer AI (22.0%), both of which were the two use cases with lowest development acceptability.
+Through further analysis of the influence of rationale factors on judgments using a mixed-effects model (Table 5), we found that use of cost-benefit reasoning in rationale is positively associated with development acceptability (), while rule-based reasoning is negatively associated ().
+Interestingly, for usage, cost-benefit reasoning was negatively associated with acceptability () but positively for judgment weighted by confidence ().
+This suggests that, although cost-benefit considerations may reduce usage acceptance, they boost confidence when judgments are favorable.
+The significant negative association of rule-based reasoning with judgments indicate that for certain contexts, the inherent action of using AI is viewed negatively.
+Participants cited diverse concerns related to AI itself (e.g., “because it would not have human sympathy”, P16), human needs (e.g., “Humans need human interactions in order to learn properly”, P44), societal impact (e.g., “Overreliance on AI …”, P132), and morality (e.g., “Having artificial intelligence try and sell you things is immoral”, P13).
+These results suggest that AI’s acceptability heavily lies in its positive outcomes but can be outweighed by established rules.
+Moral Foundations
 Beyond reasoning types, we explored moral foundations to provide insights into what values are relevant for AI use case decisions (Figure 5).
 For example, P12 responded that Elementary School Teacher AI use case should be developed because it “could give elementary schooling to children who are bed ridden…”, which was annotated with both values of Care (focusing on the well-being and nurturing of bed-ridden children) and Fairness (focusing on fair access to education).
-Upon analysis, we observe that Care (i.e., dislike of pain of others, feelings of empathy and compassion toward others) was the most prevalent moral value in participants’ rationales across the use cases.
-Of note, Care could impact the acceptability of AI use cases in both directions, as conveyed by P385 who noted that Customized Lifestyle Coach AI should be developed because “it may help improve some people’s health” but would change their decision if “it caused harm to even one person.” While our choice of medical domain in personal usage could have impacted the distribution, care was still the most prevalent considering only professional use cases.
-Participants’ rationales evoked fairness value more than care for two use cases, Lawyer AI (0.62) and Government Eligibility Interviewer (0.74) respectively.
-These results could be due to the characteristics of the use cases, such as their main purpose and function, as noted by P88, Government Eligibility Interviewer should be developed because “it might be less biased and therefore more fair in it decisions (sic)”.
-Authority was most apparent in participant rationales for Lawyer AI (0.37) and Purity for Flavorful Swaps (0.35).
-5.3.3. Switching Conditions
-We have thus far explored firmness of the decision through levels of confidence provided by users.
-We further explore the flexibility of participants’ judgments to understand possible mitigation of disagreements through concerns expressed in conditions under which they would switch their decisions (Figure 7).
-Overall, Functionality (53.3%; e.g., Medical Adivce AI should not be developed if it “consistently or had a high percentage of failure to diagnose correctly.”, P373) was the most commonly noted concern that would switch participants decision.
-This was followed by Societal Impact (42.3%; e.g., Lawyer AI should not be developed if “it puts too many human lawyers out of work”, P97), Usage (31.8%; e.g., Government Eligibility Interviewer should be developed if “it was only used to read and screen applications but not for making decisions”), Not Applicable (0.02%; e.g., will not change decision).
-Interestingly, Societal Impact (50.6%) was more prevalent followed by Functionality (46.4%) for professional use cases whereas Functionality (54.4%) was the most frequently mentioned concern for personal use cases.
-One of the reasons for frequent mention of Societal Impact in professional use cases could be due to concerns of labor replacement for professional use cases: as described by P296, if Elementary School Teacher AI “was to replace teachers with the ai to save money”, they would switch decision from should be developed to should not be developed.
-Moreover, we observe that Societal Impact was the most prevalent concern for those (31%) who thought the use case was unacceptable across all use cases and was especially prevalent for Elementary School Teacher AI (0.47).
-These diverging results show the importance of understanding granular concerns by use cases to effectively address the relevant issues.
-Participants’ responses that were identified as being both rule-based and concerned about societal impact expressed rules such as necessity of humanness as expressed by P22, “a human touch is 100% necessary” when discussing Elementary School AI, whether it be due to belief that “AI would lack empathy” (P5) or that “humans need human interactions” (P44).
-| DEV ( (SE)) | USAGE ( (SE)) | |||||
-| Rationale Factors | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
-| Reasoning Type | ||||||
-| Cost-benefit | ||||||
-| Rule-based | ||||||
-| Moral Value | ||||||
-| Care | ||||||
-| Fairness | ||||||
-| Purity | ||||||
-| Authority | ||||||
-| Switching Condition | ||||||
-| Functionality | ||||||
-| Usage | ||||||
-| Societal Impact | ||||||
-5.3.4. Rationale Factors and Judgment
-The results of analyzing the influence of rationale factors on judgments of acceptability indicated several significant effects (Table 5).
-We first verified that the factors had less than 0.5 correlation with each other.
-For acceptability of development, use of Cost-benefit reasoning in rationale was associated with increased acceptability (), while Rule-based reasoning was negatively associated with acceptability ().
-Interestingly, for usage, Cost-benefit reasoning was negatively associated with acceptability () but positively for judgment weighted by confidence ().
-Rule-based reasoning indicated lower usage acceptability for judgment weighted by confidence, consistent with acceptability of development.
-In the moral value category, Fairness in rationales had positive associations with acceptance (, ; judgment weighted by confidence).
-Analyzing switching conditions showed that participants who answered negatively to the development of use cases mentioned the concerns about Functionality () and Societal Impact to switch their decisions ().
-However, those who were positive towards development of use case indicated emphasis on Usage () as a condition to reverse their decisions.
-Thus, our findings highlight diverse perspectives on requirements and concerns of AI use cases, especially with varying perception of acceptability.
-6. Conclusion and Discussion
+Upon analysis, we observe that Care (i.e., dislike of pain of others, feelings of empathy and compassion toward others) was the most prevalent moral foundation in participants’ rationales across the use cases in both categories (48%).
+Interestingly, Care could be invoked in both positive and negative regards for AI, as conveyed by P385 who noted that Customized Lifestyle Coach AI should be developed because “it may help improve some people’s health” but would change their decision if “it caused harm to even one person.”
+Although Care was the most dominant moral foundation overall, Fairness emerged more prominently in context-specific evaluations such as Lawyer AI (41%) and Government Eligibility Interviewer AI (66%).
+These results could be due to the characteristics of the use cases, such as their main purpose and function; as noted by P88, Government Eligibility Interviewer should be developed because “it might be less biased and therefore more fair in it decisions (sic)”.
+Authority was most apparent in participant rationales for Lawyer AI (28%) and Purity for Flavorful Swaps (27%).
+Moreover, Fairness in rationales had positive associations with acceptance (, ; judgment weighted by confidence; see Table 5).
+Switching Conditions
+We further explored the flexibility of participants’ judgments to understand possible mitigation of disagreements through criteria for switching their decisions (Figure 6).
+Functionality (53%; e.g., Medical Adivce AI should not be developed if it “consistently or had a high percentage of failure to diagnose correctly.”, P373) was the most commonly noted condition for switching decisions in both directions (positive to negative and vice versa).
+This was followed by Usage (40%; e.g., Government Eligibility Interviewer should be developed if “it was only used to read and screen applications but not for making decisions”), Societal Impact (36%; e.g., Lawyer AI should not be developed if “it puts too many human lawyers out of work”, P97), and Not Applicable (7%; e.g., will not change decision).
+Interestingly, for labor-replacement use cases Societal Impact (45%) was more frequently mentioned as switching conditions followed by Functionality (43%), whereas Functionality (53%) was more frequently mentioned than Societal Impact (37%) for personal use cases.
+Frequency of Societal Impact in labor-replacement use cases could be closely linked to concerns of labor replacement: as described by P296, if Elementary School Teacher AI “was to replace teachers with the ai to save money”, they would switch their decision from positive to negative.
+Moreover, for all use cases except Elementary School Teacher AI, participants tended to switch from positive to negative decisions for lack of functionality reasons.
+In contrast, for Elementary School Teacher AI, the most common shift towards acceptability when the use case showed a positive societal impact (38%).
+As shown in Table 5, mentions of Societal Impact as conditions to switch decisions were more negatively associated with judgments (, ; judgment confidence).
+However, emphasis on Usage () as a condition to reverse their decisions was positively associated.
+RQ3 Takeaways
+Reasoning types varied by context, with rule-based reasoning more common in contested use cases and negatively associated with acceptability, while cost-benefit reasoning showed a positive association.
+The moral foundation of Care was especially salient, highlighting its importance in AI judgments.
+When explaining what might change their decisions, participants most frequently cited Functionality for personal health use cases and Societal Impact for labor-replacement, underscoring the context-dependent nature of these concerns, especially to be considered when mitigating disagreements.
+6 Conclusion and Discussion
 We conducted a study to understand how and why laypeople perceive various AI use cases as acceptable or not.
 To achieve this, we developed a survey that gathered judgments and reasoning processes from 197 participants who were demographically diverse and had varying levels of experience with AI.
-Participants were asked to provide their judgments on the acceptability of AI use cases, along with rationales for their decisions (e.g., ”Should / Should not be developed, because…”) and conditions that might change their decisions (e.g., ”I would switch my decision if…”).
+Participants were asked to provide their judgments on the acceptability of AI use cases, along with rationales for their decisions (e.g., “Should / Should not be developed, because…”) and conditions that might change their decisions (e.g., “I would switch my decision if…”).
 The survey covered ten different AI use cases, spanning both personal and professional domains, and included varying levels of risk.
@@
 Professional use cases displayed more variability, notably with Elementary School Teacher AI, which was uniquely unacceptable.
-This underscores the necessity for further research into how AI should be developed and integrated, as well as what skills it should have, particularly in fields where empathy and care are crucial (Wu et al., 2024; Kawakami et al., 2024; Borg and Read, 2024).
-In addition, prior FAccT research have also highlighted how AI practitioners desire understanding lay people’s perception on AI fairness in specific use cases (Sonboli et al., 2021; Deng et al., 2022; Smith et al., 2023; Deng et al., 2023).
-Drawing from prior HCI and AI research (Deng et al., 2025; Lee et al., 2019; Cheng et al., 2019), future FAccT researchers and practitioners should explore how to meaningfully connect lay people’s use case perceptions with AI developers’ workflows.
-While prior research has emphasized understanding AI consequences (Kieslich et al., 2024) and providing tools and processes to uncover impact (Wang et al., 2024; Buçinca et al., 2023; Deng et al., 2024), our findings reveal a greater presence of rule-based reasoning in contentious use cases, suggesting a need for diverse approaches to understanding AI beyond mere consequence anticipation.
+This underscores the necessity for further research into how AI should be developed and integrated, as well as what skills it should have, particularly in fields where empathy and care are crucial (Wu et al.
+2024; Kawakami et al.
+2024; Borg and Read 2024).
+In addition, prior research have also highlighted how AI practitioners desire understanding lay people’s perception on AI fairness in specific use cases (Sonboli et al.
+2021; Deng et al.
+2022; Smith, Beattie, and Cramer 2023; Deng et al.
+2023).
+Drawing from prior HCI and AI research (Deng et al.
+2025; Lee et al.
+2019; Cheng et al.
+2019), future researchers and practitioners should explore how to meaningfully connect lay people’s use case perceptions with AI developers’ workflows.
+While prior research has emphasized understanding AI consequences (Kieslich, Helberger, and Diakopoulos 2024) and providing tools and processes to uncover impact (Wang et al.
+2024; Buçinca et al.
+2023; Deng, Barocas, and Vaughan 2024), our findings reveal a greater presence of rule-based reasoning in contentious use cases, suggesting a need for diverse approaches to understanding AI beyond mere consequence anticipation.
 Moreover, while care was generally predominant, we observed that fairness gained prominence in Lawyer AI and Government Eligibility Interviewer AI.
-This variability underscores the importance of considering values in AI evaluation and training (Barocas et al., 2021; Bhardwaj et al., 2024), rather than solely emphasizing functionality, which is the current trend in AI research (Birhane et al., 2022).
-Additionally, societal impact considerations were more evident in unacceptable use cases, emphasizing the necessity for implementing safety guardrails when deploying AI with significant social implications (Solaiman, 2023).
+This variability underscores the importance of considering values in AI evaluation and training (Barocas et al.
+2021; Bhardwaj et al.
+2024), rather than solely emphasizing functionality, which is the current trend in AI research (Birhane et al.
+2022).
+Additionally, societal impact considerations were more evident in unacceptable use cases, emphasizing the necessity for implementing safety guardrails when deploying AI with significant social implications (Solaiman 2023).
 Demographics and AI Literacy
-In line with prior work (Kingsley et al., 2024; Mun et al., 2024), our results highlighted significant differences among demographic groups and perceived acceptance of use cases, especially for professional use (§5.2).
+In line with prior work (Kingsley et al.
+2024; Mun et al.
+2024), our results highlighted significant differences among demographic groups and perceived acceptance of use cases, especially for professional use (§5.2).
 Non-majority demographic groups, especially non-male gender groups, found both personal and professional use cases less acceptable.
 Those experiencing high discrimination chronicity also found professional use less acceptable.
-Our findings provide future FAccT research with valuable empirical insights on AI integration in contexts such as workplace, where marginalized worker’s agency, earning, and occupational well-being are disproportionately affected (Ming et al., 2024; Alcover et al., 2021).
+Our findings offer empirical insights for future research on AI integration in workplaces, where marginalized workers’ agency, income, and well-being are disproportionately impacted (Ming et al.
+2024; Alcover et al.
+2021).
 Furthermore, our work highlighted a potential polarization on perceptions of AI among workers as those with 40+ hours employment and those who had advanced degrees were more positive towards AI use cases, suggesting that the relationship stakeholders have to AI and jobs might influence acceptability.
-Such a concern was expressed by one of our participants who opposed development of Telemarketer AI because it “overlaps with my industry, and hence serves as a threat to my job security” (P35).
-Thus, our results corroborate the need to further explore methods to include diverse workers and various stakeholders into the discussion of workplace AI integration and development (Fox et al., 2020; Cheon, 2023).
+This concern was expressed by a participant who opposed the development of Telemarketer AI, stating that it “overlaps with my industry, and hence serves as a threat to my job security” (P35).
+Thus, our results corroborate the need to further explore methods to include diverse workers and various stakeholders into the discussion of workplace AI integration and development (Fox et al.
+2020; Cheon 2023).
 We also found that frequent AI usage increased acceptance, while understanding AI ethics and limitations decreased acceptance.
-This suggests that balanced AI awareness and education, encompassing usage, skills, and ethics, could guide and improve decision-making (Raji et al., 2021), e.g., through educational interventions targeting AI skills and ethical implication literacy (e.g., (Wong and Nguyen, 2021; Shen et al., 2021)).
+This suggests that balanced AI awareness and education, encompassing usage, skills, and ethics, could guide and improve decision-making (Raji, Scheuerman, and Amironesei 2021), e.g., through educational interventions targeting AI skills and ethical implication literacy (e.g., Wong and Nguyen 2021; Shen et al.
+2021).
 Rationales
-Through analyzing participants’ rationales, we observed an interesting pattern with higher cost-benefit reasoning use cases with least disagreement and more prevalence of rule-based reasoning for those with higher disagreement (§5.3).
-These results suggest different valuation systems employed by participants perhaps leading to different conclusions and suggests some use cases have beyond simple utilitarian implication for society, which should be explored more carefully in future studies.
-Building upon our empirical findings, future work could develop tools and interventions which encourage specific types of acceptability reasoning such as rule and value based (Sorensen et al., 2024) or cost-benefit analyses (Li et al., 2024).
+Through analyzing participants’ rationales, we observed an interesting pattern: use cases with less disagreement tended to elicit more cost-benefit (utilitarian) reasoning, while those with greater disagreement showed more rule-based (deontological) reasoning (§5.3).
+This suggests participants may apply different valuation frameworks, leading to diverging judgments, and highlights that some use cases raise concerns beyond simple utilitarian considerations.
+Our results thus underscore crucial elements of participants decision making pattern when only assessing impact as many prior works have done.
+Building upon our empirical findings, future work could develop diverse open-ended analysis for eliciting deliberations as well as tools and interventions using specific types of acceptability reasoning such as rule and value based (Sorensen et al.
+2024) or cost-benefit analyses (Li et al.
+2024).
 However, as our study was limited to the two reasoning type categories, expanding this analysis would be essential for future work including finding ways to classify what features people are considering in their decisions, how the weights on those features impact what kind of decision-making strategy they will use, and whether there are other ways to understand their decision strategies beyond our current classification.
-Future FAccT research can build upon these further understandings to guide policy making and consensus building.
-For example, in addition to conducting surveys with single participant, future work can explore how group discussions and deliberations shape communities’ collective understanding of AI impacts (e.g., (Kuo et al., 2024; Lee et al., 2019; DeVos et al., 2022; Gordon et al., 2022; Zhang et al., 2023))
+Future research can build upon these further understandings to guide policy making and consensus building.
+For example, future work could explore how group discussions, beyond individual surveys, shape communities’ collective understanding of AI impacts (e.g., Kuo et al.
+2024; Lee et al.
+2019; DeVos et al.
+2022; Gordon et al.
+2022; Zhang et al.
+2023).
 Appendix A Additional Survey Details
-A.1. Survey Questions
+A.1 Survey Questions
 In our survey we ask participants to read one or more descriptions of AI use cases and to make two judgments: 1) “Do you think a technology like this should exist?” (Q1) and 2) “If the ¡use case¿ exists, would you use its services?” (Q5).
@@
 | Question ID | Question |
-Moral Foundation Questionnaire (First Half)
-When you decide whether something is right or wrong, to what extent is the following consideration relevant to your thinking? |
+| Moral Foundation Questionnaire (First Half) | |
+| When you decide whether something is right or wrong, to what extent is the following consideration relevant to your thinking? | |
 | MFQ 1 | Whether or not someone suffered emotionally |
@@
 | MFQ 22 | I would call some acts wrong on the grounds that they are unnatural. |
-A.2. Participant Details
-The demographics of the participants for our study is shown in Tables 13 to 35.
+A.2 Use Case Descriptions
+See Table12 for labor replacement use cases as shown to the participants and Table13 for personal health use cases.
+| Use Case | Education Level | Description |
+| Lawyer AI | High: Doctoral or Professional Degree, Master’s Degree | Imagine an AI system that performs the tasks of a Lawyer digitally.
+The AI system advises and represents clients on digital legal proceedings or transactions.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Lawyer AI in the rest of the survey.
+|
+| Elementary School Teacher AI | Middle High: Bachelor Degree, Associate Degree | Imagine an AI system that performs the tasks of an Elementary School Teacher digitally.
+The AI system teaches academic skills to students at the elementary school level through online interactions.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Elementary School Teacher AI in the rest of the survey.
+|
+| IT Support Specialist AI | Middle: Some College No Degree, Postsecondary Nondegree Award | Imagine an AI system that performs the tasks of an IT Support Specialist digitally.
+The AI system maintains computer networks and provides technical help to computer users.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as IT Support Specialist AI in the rest of the survey.
+|
+| Government Eligibility Interviewer AI | Middle Low: High School Diploma or Equivalent | Imagine an AI system that performs the tasks of an Eligibility Interviewer for Government Programs digitally.
+This AI system determines eligibility of persons applying to receive assistance from government programs and agency resources, such as welfare, unemployment benefits, social security, and public housing.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Government Eligibility Interviewer AI in the rest of the survey.
+|
+| Telemarketer AI | Low: No Formal Educational Credential | Imagine an AI system that performs the tasks of a Telemarketer digitally.
+The AI system solicits donations or orders for goods or services over the telephone.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Telemarketer AI in the rest of the survey.
+|
+| Use Case | EU Risk Level | Description |
+| Digital Medical Advice AI | High | Imagine an AI system that provides preliminary medical assessments to help patients get efficient medical consultations and treatment (i.e., medical advice).
+It analyzes physical characteristics, reported symptoms, and medical history to suggest potential health issues, provide treatment plans, or direct patients to appropriate medical professionals and facilities.
+In emergencies, it can also act as a triage tool used by medical institutions to prioritize care.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Digital Medical Advice AI in the rest of the survey.
+|
+| Customized Lifestyle Coach AI | High / Limited | Imagine an AI system that offers personalized advice on how to manage healthy lifestyles and enhance wellness (i.e., customized lifestyle coaching).
+Beyond health tracking, it provides actionable insights on integrating fitness routines, dietary adjustments, and mindfulness practices into your daily schedule.
+The health data it collects can also be used to provide accurate information for life and health insurance.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Customized Lifestyle Coach AI in the rest of the survey.
+|
+| Personal Health Research AI | Limited | Imagine an AI system that summarizes and documents complex research findings related to personal health issues users are interested in (i.e., personalized health research findings).
+It translates medical jargon into accessible language, explains the latest studies relevant to conditions of user interest, and provides a summary of the overall findings.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Personal Health Research AI in the rest of the survey.
+|
+| Nutrition Optimizer AI | Limited / Low | Imagine an AI system that organizes meal schedules and optimizes nutritional intake based on a user’s specific health goals (i.e., optimal nutrition plan).
+It crafts recipes, meal plans, and grocery lists optimized for users’ culinary preferences, dietary needs, fitness ambitions, and lifestyle.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Nutrition Optimizer AI in the rest of the survey.
+|
+| Flavorful Swaps AI | Low | Imagine an AI system that aims to enhance users’ diets by suggesting delicious, health-conscious alternatives to unhealthy favorite foods (i.e., healthy flavorful swaps).
+It offers meal swaps that maintain satisfying flavors while being aligned with users’ health and dietary goals.
+The system can read, write, and talk fluently.
+It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data.
+We will refer to the AI system as Flavorful Swaps AI in the rest of the survey.
+|
+A.3 Participant Details
+The demographics of the participants for our study is shown in Tables 15 to 37.
 There was a fairly balanced distribution of participants across the different age groups, although there was a slightly higher proportion of participants in the 25-34 years old and 45-54 years old age ranges.
@@
 | 7 | 45 | 85 | 65 | — | — |
-| Score | AI Ethics |
+| AI Ethics Score | Count |
 | 5 | 15 |
@@
 | 18 | 35 |
-A.3. Open-text Annotation Dimensions
+A.4 Open-text Annotation Dimensions
 Reasoning Type
-Inspired by previous works in moral psychology, we used two main reasoning types to characterize participants’ decision making pattern as expressed in their open-text answers: cost-benefit reasoning and rule-based reasoning (Cheung et al., 2024).
+Inspired by previous works in moral psychology, we used two main reasoning types to characterize participants’ decision making pattern as expressed in their open-text answers: cost-benefit reasoning and rule-based reasoning (Cheung, Maier, and Lieder 2024).
 These two reasoning types are rooted in two decision making processes in moral and wider decision making literature: utilitarian and deontological reasoning, respectively.
@@
 Moral Values
-To annotate which values were prevalent in participants’ considerations of use cases, we used five moral values: care, fairness, loyalty, authority, and purity (Graham et al., 2011, 2008).
-While these dimensions have been re-defined to include more diverse values from participants beyond WEIRD (white, educated, industrialized, rich, and democratic) (Atari et al., 2023), we used these five dimensions due to brevity of the questionnaire, which was used in the survey to provide importance of each values to participants.
+To annotate which values were prevalent in participants’ considerations of use cases, we used five moral values: care, fairness, loyalty, authority, and purity (Graham et al.
+2011, 2008).
+While these dimensions have been re-defined to include more diverse values from participants beyond WEIRD (white, educated, industrialized, rich, and democratic) (Atari et al.
+2023), we used these five dimensions due to brevity of the questionnaire, which was used in the survey to provide importance of each values to participants.
 Switching Conditions
@@
 and user concern annotation practice adopted by Mun et al..
-Appendix B Open-text Annotation Details
-B.1. Automatic Annotation
-B.1.1. Methods
-We used Open-AI’s o1-mini model with maximum tokens set to 1024 to control response length, use a temperature of 0.7 to manage randomness, and keep top_p at 1 with default settings for frequency and presence penalties at 0.
+Appendix B Extended Results
+Figure 7 shows the distribution of participants’ judgment variation across categories as discussed in §5.1.
+Appendix C Open-text Annotation Details
+C.1 Automatic Annotation
+Methods
+We used Open-AI’s gpt-4o model with maximum tokens set to 1024 to control response length, use a temperature of 0.7 to manage randomness, and keep top_p at 1 with default settings for frequency and presence penalties at 0.
 Prompts will be released with data upon acceptance.
-B.1.2. Results
-Results for inter-rater reliability analysis of o1’s annotations are shown in Table 21.
+Results
+Results for inter-rater reliability between human and gpt-4o annotations are shown in Table 23.
 | Category | AC1 | Interpretation | 95% CI | p-value | z | SE | PA | PE |
@@
 | Rule Based | 0.848 | Almost Perfect | (0.754, 0.943) | 0.0 | 17.8 | 0.0476 | 0.890 | 0.276 |
-| Care | 0.427 | Moderate | (0.234, 0.619) | 4.40 | 0.0970 | 0.650 | 0.390 | |
-| Fairness | 0.411 | Moderate | (0.224, 0.599) | 4.35 | 0.0945 | 0.690 | 0.474 | |
-| Authority | 0.839 | Almost Perfect | (0.743, 0.935) | 0.0 | 17.3 | 0.0486 | 0.880 | 0.255 |
-| Purity | 0.758 | Substantial | (0.639, 0.878) | 0.0 | 12.6 | 0.0603 | 0.820 | 0.255 |
-| Functionality | 0.587 | Moderate | (0.425, 0.749) | 7.18 | 0.0818 | 0.790 | 0.492 | |
-| Usage | 0.673 | Substantial | (0.525, 0.822) | 9.03 | 0.0746 | 0.820 | 0.449 | |
-| Societal Impact | 0.595 | Moderate | (0.432, 0.759) | 7.23 | 0.0823 | 0.770 | 0.432 |
-Appendix C Factors Impacting Acceptability Judgments
-C.1. Use Case Factors
+| Care | 0.705 | Substantial | (0.563, 0.847) | 9.88 | 0.0713 | 0.850 | 0.492 | |
+| Fairness | 0.609 | Substantial | (0.448, 0.769) | 7.53 | 0.0808 | 0.790 | 0.464 | |
+| Authority | 0.897 | Almost Perfect | (0.822, 0.972) | 0.0 | 23.70 | 0.0378 | 0.920 | 0.226 |
+| Purity | 0.801 | Almost Perfect | (0.693, 0.908) | 0.0 | 14.77 | 0.0542 | 0.850 | 0.248 |
+| Functionality | 0.701 | Substantial | (0.559, 0.844) | 9.79 | 0.0716 | 0.850 | 0.498 | |
+| Usage | 0.674 | Substantial | (0.527, 0.822) | 9.05 | 0.0745 | 0.830 | 0.478 | |
+| Societal Impact | 0.566 | Moderate | (0.397, 0.735) | 6.66 | 0.0851 | 0.750 | 0.424 |
+Appendix D Factors Impacting Acceptability Judgments
+D.1 Use Case Factors
 Additional analysis of use case factors showing distribution of judgments by use case sorted by standard deviation is shown in Figure 8. Table 24 shows analysis of use case effect using ANOVA.
-| Acceptability | Aspect | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
+| Acceptability | Aspect | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(F) |
 | EXIST | Judgment | Category | 29.903 | 29.903 | 1 | 197 | 5.98e-11 *** |
-| Use Case | 86.116 | 9.5684 | 9 | 641.38 | ¡ 2.2e-16 *** | ||
+| Use Case | 86.116 | 9.5684 | 9 | 641.38 | 2.2e-16 *** | ||
 | Confidence | Category | 0.0017113 | 0.0017113 | 1 | 197 | 0.9563 | |
@@
 | USAGE | Judgment | Category | 8.4257 | 8.4257 | 1 | 197 | 0.0002488 *** |
-| Use Case | 73.801 | 8.2001 | 9 | 610.34 | ¡ 2.2e-16 *** | ||
+| Use Case | 73.801 | 8.2001 | 9 | 610.34 | 2.2e-16 *** | ||
 | Confidence | Category | 1.3444 | 1.3444 | 1 | 197 | 0.153 | |
 | Use Case | 20.721 | 2.3023 | 9 | 603.36 | 0.0001783 *** |
-C.2. Demographics Factors
+D.2 Demographics Factors
 Additional analysis using ANOVA for demographic factors is shown in Table 25.
@@
 | Var: Residual | ||||||
-C.2.1. Questionnaires
+Questionnaires
 Interestingly, only Loyalty had a significant effect on both existence () and usage () as shown in Table 26.
@@
 However, Loyalty, as shown in Figure 5, does not appear as frequently in participants’ open text responses compared to values such as Care and Fairness and is the only value that did not have a significant association with use cases.
-Appendix D Factors in Participant Rationale
-D.1. Reasoning Types
+Appendix E Factors in Participant Rationale
+E.1 Reasoning Types
 We show the flow of participants’ decisions and corresponding rationales throughout use cases in Figure 9, which shows interesting distribution and switching of reasoning types, which would be interesting for future studies to consider.
@@
 | Var: Residual | ||
-D.2. Impact of Rationale Factors on Judgment
+E.2 Impact of Rationale Factors on Judgment
 We display the analysis results using ANOVA to understand the effect of rationale factors on judgment in Table 28.
-| Acceptability | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
+| Acceptability | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(F) |
 | EXIST | Judgment | |||||
-| Cost Benefit | 2.941 | 2.941 | 1 | 955.32 | 0.001306 ** | |
-| Rule Based | 7.187 | 7.187 | 1 | 941.74 | 5.562e-07 *** | |
-| Fairness | 1.551 | 1.551 | 1 | 966.41 | 0.019414 * | |
-| Authority | 1.044 | 1.044 | 1 | 961.50 | 0.055021 . | |
-| Functionality | 1.137 | 1.137 | 1 | 964.09 | 0.045230 * | |
-| Usage | 125.594 | 125.594 | 1 | 837.10 | ¡ 2.2e-16 *** | |
-| Societal Impact | 1.873 | 1.873 | 1 | 967.89 | 0.010228 * | |
+| Cost Benefit | 2.764 | 2.764 | 1 | 953.92 | 0.0017 ** | |
+| Rule Based | 6.939 | 6.939 | 1 | 937.59 | 7.57e-07 *** | |
+| Fairness | 3.073 | 3.073 | 1 | 969.63 | 0.00095 *** | |
+| Authority | 3.941 | 3.941 | 1 | 973.72 | 0.00018 *** | |
+| Usage | 127.631 | 127.631 | 1 | 828.54 | 2.2e-16 *** | |
 | Confidence | ||||||
-| Rule Based | 5.7805 | 5.7805 | 1 | 856.38 | 0.001158 ** | |
-| Care | 3.8723 | 3.8723 | 1 | 907.71 | 0.007762 ** | |
-| Fairness | 5.7117 | 5.7117 | 1 | 899.21 | 0.001237 ** | |
-| Authority | 2.2665 | 2.2665 | 1 | 851.91 | 0.041526 * | |
-| Usage | 2.6254 | 2.6254 | 1 | 809.17 | 0.028303 * | |
+| Rule Based | 5.5845 | 5.5845 | 1 | 854.68 | 0.00149 ** | |
+| Fairness | 3.9022 | 3.9022 | 1 | 893.72 | 0.00787 ** | |
 | Judgment x Confidence | ||||||
-| Cost Benefit | 48.04 | 48.04 | 1 | 940.91 | 0.0004985 *** | |
-| Rule Based | 86.12 | 86.12 | 1 | 924.15 | 3.331e-06 *** | |
-| Fairness | 31.87 | 31.87 | 1 | 972.26 | 0.0045236 ** | |
-| Authority | 14.24 | 14.24 | 1 | 972.97 | 0.0574858 . | |
-| Functionality | 18.99 | 18.99 | 1 | 971.66 | 0.0283005 * | |
-| Usage | 2454.30 | 2454.30 | 1 | 888.43 | ¡ 2.2e-16 *** | |
-| Societal Impact | 34.26 | 34.26 | 1 | 971.98 | 0.0032486 ** | |
+| Cost Benefit | 44.72 | 44.72 | 1 | 936.85 | 0.00072 *** | |
+| Rule Based | 82.21 | 82.21 | 1 | 917.67 | 4.79e-06 *** | |
+| Fairness | 43.81 | 43.81 | 1 | 971.97 | 0.00081 *** | |
+| Authority | 56.43 | 56.43 | 1 | 965.79 | 0.00015 *** | |
+| Usage | 2458.62 | 2458.62 | 1 | 887.02 | 2.2e-16 *** | |
+| Societal Impact | 22.15 | 22.15 | 1 | 969.86 | 0.0171 * | |
 | USAGE | Judgment | |||||
-| Cost Benefit | 0.28 | 0.28 | 1 | 933.91 | 0.01742 * | |
-| Usage | 491.35 | 491.35 | 1 | 943.16 | ¡ 2e-16 *** | |
+| Cost Benefit | 0.28 | 0.28 | 1 | 930.56 | 0.0177 * | |
+| Usage | 496.98 | 496.98 | 1 | 937.06 | 2e-16 *** | |
 | Confidence | ||||||
-| Cost Benefit | 2.2380 | 2.2380 | 1 | 864.41 | 0.0555024 . | |
-| Rule Based | 2.0230 | 2.0230 | 1 | 852.41 | 0.0686374 . | |
-| Fairness | 4.2517 | 4.2517 | 1 | 895.25 | 0.0083622 ** | |
-| Authority | 2.4525 | 2.4525 | 1 | 852.20 | 0.0450309 * | |
-| Loyalty | 2.2751 | 2.2751 | 1 | 875.10 | 0.0535163 . | |
-| Functionality | 2.6366 | 2.6366 | 1 | 897.34 | 0.0376892 * | |
-| Usage | 9.2673 | 9.2673 | 1 | 817.52 | 0.0001033 *** | |
-| Societal Impact | 2.1237 | 2.1237 | 1 | 913.41 | 0.0620930 . | |
+| Cost Benefit | 2.4825 | 2.4825 | 1 | 863.46 | 0.0443 * | |
+| Care | 3.6966 | 3.6966 | 1 | 812.12 | 0.0142 * | |
+| Usage | 11.9628 | 11.9628 | 1 | 813.76 | 1.12e-05 *** | |
 | Judgment x Confidence | ||||||
-| Cost Benefit | 106.358 | 106.358 | 1 | 874.26 | 8.333e-05 *** | |
-| Rule Based | 33.512 | 33.512 | 1 | 859.58 | 0.026742 * | |
-| Fairness | 89.120 | 89.120 | 1 | 930.00 | 0.000312 *** | |
-| Societal Impact | 74.975 | 74.975 | 1 | 929.00 | 0.000938 *** | |
-D.3. Factors Influencing Moral Foundations in Rationale
+| Cost Benefit | 143.66 | 143.66 | 1 | 872.38 | 4.92e-06 *** | |
+| Fairness | 72.485 | 72.485 | 1 | 918.88 | 0.00113 ** | |
+| Societal Impact | 69.555 | 69.555 | 1 | 932.22 | 0.00143 ** | |
+E.3 Factors Influencing Moral Foundations in Rationale
 In Table 29 we display analysis result using linear mixed effects on factors that may influence moral foundations appealed to in participants’ rationales.
@@
 | Var: prolific_id (Intercept) | |||||
-Appendix E Survey 2: Explicit Weighing of Harms and Benefits of Use Cases
+Appendix F Survey 2: Explicit Weighing of Harms and Benefits of Use Cases
 Although not included in main text, we administered a variation of our main study where we asked participants to explicitly reason through harms and benefits.
@@
 However, we saw almost no effect.
-E.1. Study Overview
+F.1 Study Overview
 To better understand the reasoning behind participants decisions about the judgment and usage of the use cases, we conducted a second study with 1 survey for each category (Labor Replacement Use Cases and Personal Use Cases).
@@
 Furthermore, we asked participants the judgment and usage decision questions before and after the set of harms and benefits questions to see if listing out reasons about an use case would elicit any change in their decisions.
-The details of the second study can be found in E.2 and the results can be found in E.3.
-E.2. Setup and Details
+The details of the second study can be found in F.2 and the results can be found in F.3.
+F.2 Setup and Details
 While these same set of questions are asked for all five use cases for our main study, in our second study, participants are randomly allocated a single use case.
@@
 To understand how the judgment and usage decisions are affected by other factors, we asked the participants about their demographics, ai literacy levels and several other reasoning factors after the main set of questions, and these questions can be found in §A.1 The main questions for the second study can be found in Table 31.
-The participant demographics for the second study can be found in Tables 30 to 35.
+The participant demographics for the second study can be found in Tables 32 to 37.
 The distribution of each use case within each scenario (Labor Replacement Use Cases and Personal Use Cases) for the second study is relatively well-balanced and can be found in Table 30.
@@
 | Accommodation and Food Services | 4 (4.1) | Jewish | 1 (1.0) |
-E.3. Results
-To explore the possible impact of explicitly weighing harms and benefits of a use case on participant’s decision, we analyzed the participant’s judgment of acceptability before and after explicit weighing of harms and benefits (Study 2; see § LABEL:sssec:survey-qs for details on questions asked).
+F.3 Results
+To explore the possible impact of explicitly weighing harms and benefits of a use case on participant’s decision, we analyzed the participant’s judgment of acceptability before and after explicit weighing of harms and benefits (Study 2; see Table31 for details on questions asked).
 The Type III ANOVA with Satterthwaite’s method for measurement time (before, after) indicated a marginally significant effect on usage judgment weighed by confidence, which suggests that explicit harms and benefits weighing may have an influence, albeit not at conventional significance levels.